### v1

# Improving Technical ”How-to” Query Accuracy with Automated Search Results Verification and Reranking

Abstract.
Many people use search engines to find online guidance to solve computer or mobile device problems. Users frequently encounter challenges in identifying effective solutions from search results, often wasting time trying ineffective solutions that seem relevant yet fail to solve the real problems. This paper introduces a novel approach to improving the accuracy and relevance of online technical support search results through automated search results verification and reranking. Taking ”How-to” queries specific to on-device execution as a starting point, we first developed a solution that allows an AI agent to interpret and execute step-by-step instructions in the search results in a controlled Android environment. We further integrated the agent’s findings into a reranking mechanism that orders search results based on the success indicators of the tested solutions.
The paper details the architecture of our solution and a comprehensive evaluation of the system through a series of tests across various application domains. The results demonstrate a significant improvement in the quality and reliability of the top-ranked results. Our findings suggest a paradigm shift in how search engine ranking for online technical support help can be optimized, offering a scalable and automated solution to the pervasive challenge of finding effective and reliable online help.
1. Introduction
Often people lack the knowledge to complete a task and rely on a search engine to retrieve potential instructions by asking ”How-to” or similar style queries. These requests are ”procedural queries” where the user’s intent is to retrieve step-by-step instructions for performing a specific task or operation. However, instructions extracted from retrieved results may not be executable, as the search engine primarily ranks pages based on similiarty of search query and candidate web page (Qin and Liu, 2013). Furthermore, although user engagement information, such as search logs and page quality indicators like PageRank, are utilized for ranking results in modern search engines, the optimal solution often varies on a user-specific basis. The use of different operating systems, screen types, and app versions motivates a personalized reranker. In addition, search engines integrated with large language models (LLMs) haven shown to involve the likelihood of producing hallucinations (Lewis et al., 2020).
How do users solve the problem? Typically, a user could visit multiple retrieved pages. If a page that looks relevant contains some step-by-step instructions, the user might try to follow the instructions to see whether it solves the problem. However, users frequently experience frustration and waste considerable time attempting various instructions across different pages. Can a search engine assist the user in a more efficient way? Motivated by the users’ time-consuming manual solution, we substitute this tedious process with an automatic instruction verification process as an additional component of the search process. Our hypothesis is that search results verification for technical ”How-to” queries can achieve decent accuracy given the recent research progress on multimodal LLMs, especially with the help of state-of-the-art GPT agents.
In particular, we introduce a three-stage solution to do instruction set verification and reranking relevant pages, as illustrated in Figure 1: The first stage involves an Instruction Extraction Model to obtain step-by-step instructions from each retrieved page for a ”How-to” search query. The second stage verifies their quality by automatically trying or simulating the instruction execution on devices. The third stage reranks retrieved results based on the execution information.
We have developed a comprehensive end-to-end solution to rerank ”How-to” search query results on Android systems and applications across domains. This can be enhanced as a platform-agnostic approach with slight adaptions to support other environments, such as web or desktop. Our solution enhances search engine results by reranking pages based on a simulated verification on Android devices and preliminary experimental outcomes have demonstrated the effectiveness of our methods. Our key contributions are outlined as follows:
We propose adding search result verification and reranking into the search process.
For technical ”How-to” queries, we proposed a three-stage solution. At the first stage, we propose an information extraction solution to extract step-by-step instructions from web pages using generative AI with grounding techniques. At the second stage, we propose an action agent to take instructions in natural language and execute a series of actions on client device. At the third stage, we introduce a range of features that integrate both textual data and the outcomes of executions, and these features are then utilized to rerank the retrieved results.
To illustrate and support future research in this direction, we developed a new research platform MagicWand 111http://nlp-in-477-l.soe.ucsc.edu:8000/admin/ that can verify and rerank search results for ”How-to” queries across several different domain applications on Android. We also collected a new dataset with human annotators.
We evaluated the proposed idea and the results show it significantly enhances the performance of a leading baseline search engine (i.e. Google).
2. Related Work
As far as we know, there is no prior work on search results verification for ”How-to” queries using simulated execution. On the other hand, the techniques we choose to use when developing different components of the proposed solution are inspired by various existing works from multiple domains. Therefore, we briefly review three important domains that influenced the technical choices we made in our research.
2.1. Information Extraction
Information extraction from web pages (Bevendorff et al., 2023), as a broader concept of our extraction task, was studied heavily back in 1990s (Finn et al., 2001). There are two methodologies: heuristics based approaches and machine learning based approaches. Heuristics approaches apply either a single metric or a set of rules along a hierarchy HTML tree to pinpoint the main content blocks. For instance, (Sun et al., 2011) considered ”text density” and the DOM tree structure, while (Pomikálek, 2011) segmented the HTML tree and then used heuristics for classification. In comparison, machine learning approaches in early studies involve sequence labeling methods and deep neural networks: (Vogels et al., 2018) employed a hidden Markov model and a convolutional neural network to classify web regions; (Jung et al., 2021) used browser-rendered visual features to enhance model performance.
Recently, researchers also leveraged LLMs to facilitate relevant downstream tasks. For example, (Gur et al., 2022) enabled T5-based models for semantic classification of HTML elements, description generation for HTML input, and autonomous web navigation. DOM-LM (Deng et al., 2022), MarkupLM (Li et al., 2021) used transformer-based architecture to represent the DOM tree and further improve classification accuracy. A new dataset PLAtE of list-like websites, such as online stores, was developed by Amazon research team to analyze how LLMs work on structure extraction (San et al., 2022). Researchers also endeavored to better utilize pre-trained LLMs to extract target information to avoid training LLMs. The instruction extraction approach used in our research is inspired by the prior work that used zero-shot GPT3 to extract plans from text (Hernandez et al., 2021).
2.2. Acting Agent
Enabling intelligent agents to behave according to instructions involves two main tasks: the first is empowering agents with sufficient knowledge to make decisions and act skillfully. The second is executing actions, collecting feedback, and recognizing environmental changes caused by these actions. To achieve the first, researchers often leverage and enhance LLMs’ reasoning and tooling capabilities (Mialon et al., 2023). For the second, agents should be enabled to interact with the host system to understand feasible actions and their results. Recent studies include using pre-trained GPT2 (Radford et al., 2019) in household simulations (Li et al., 2022a), RT1 (Brohan et al., 2022) and RT2 (Zitkovich et al., 2023) in robot control. This section will focus on how to adapt these tasks to the Android platform, on which we will carry out our experiments.
Acting Agent on Android:
An early work on Android was (Li et al., 2019) using LSTM to predict fixed action and parameter pairs on dataset RICO (Deka et al., 2017). Later, Seq2Act (Li et al., 2020) utilized Transformers for more flexible action execution given instructions. This approach set a strong baseline on a new dataset UGIF (Venkatesh et al., 2022), further enhanced by META-GUI (Sun et al., 2022) with a proposed multimodal Action Prediction Model for conversational scenarios. Along this work, Auto-UI (Zhan and Zhang, 2023) built an ”Action Chain” based model using UI representation extracted from BLIP-2 (Li et al., 2023) as visual input, which outperformed two other competitive approaches (i.e. Behavior Cloning and Simple action Chain of Thoughts (Wei et al., 2022)) on the AWIT dataset (Rawles et al., 2023). However, they overlooked the Android UI Control hierarchy, and the transformers trained without such knowledge can not predict actions accurately given ambiguous or sparse instructions. (Huang et al., 2022a) demonstrated that LLMs equipped with domain knowledge could break down high-level tasks into mid-level plans. Researchers were also starting to explore multi-modal LLM agents in Android automation, as evidenced by Yan et al. (2023); Yang et al. (2023b). Our Action Prediction Model aligns with this line of research, however, we’ve refined it by distilling the UI Control list from runtime metadata and prompting GPT4-V based on the context.
Action Execution:
Usually, researchers resorted to two technologies to detect UI Hierarchy and take actions on Android devices: One is Android Debug Bridge (ADB) (Guideline, 2023a), such as Toyama et al. (2021); Yang et al. (2023b), the other is Accessibility Service API (Guideline, 2023b), like Salehnamadi et al. (2022); Salehnamadi (2022). ADB-based solutions, like AndroidEnv (Toyama et al., 2021), adbutils (adb library for adb service, 2022), need additional connection settings for Android devices and the paired computer, which is unrealistic for end users. In comparison, Accessibility Service-based solutions can be delivered as standalone mobile applications due to the availability of Accessibility Service API on most systems. In our experiments, we used the Accessibility Service API.
2.3. Reranking
In information retrieval (IR), reranking is the adjustment of document rankings based on query-document relevance after the initial retrieval. The evolution of reranking began with RankNet (Burges et al., 2005) in 2005, applying gradient descent in the process of learning a ranking function. LambdaRank (Burges, 2010) later refined this approach by optimizing the gradients of the model scores and addressing issues caused by discontinuous gradients. And LambdaMART, combining MART (Friedman, 2001) with LambdaRank, achieved victory in the 2010 Yahoo! Learning-to-Rank Challenge. Subsequently, neural ranking models like DRMM (Guo et al., 2016) and Duet (Mitra et al., 2017) emerged. Recently, (Pobrotyn et al., 2020) introduced a context-aware neural network for item relevance calculation, and (Qin et al., 2020) proposed DASALC, a neural LTR model yielding performance comparable to LambdaMART. This progress inspires our adoption of NeuralNDCG (Pobrotyn and Bialobrzeski, 2021), which utilizes NeuralSort (Grover et al., 2019) for end-to-end, gradient-based stochastic optimization using a NDCG-based sorting operator.
3. Methods
The flowchart in Figure 1 outlines our proposed process to improve the accuracy and usefulness of ”How-to” search query results. It achieves this by extracting instructions from web pages and validating them through execution on the device. The validation outcomes are then used to refine and adjust the search result rankings. The system can overcome some limitations of traditional ranking algorithms by incorporating user-specific information (device, OS version, application version, etc.) and collecting verification results from client devices.
This is a three-stage process. Stage 1 is Instruction Extraction: given the initial top N retrieved results, a generative AI model processes each document to generate a sequence of instructions. Then, a grounding module analyzes the structure of the HTML web pages and further improves the quality of the extracted instructions. Stage 2 is On-device Execution: the Action Prediction Model reads the instructions extracted from stage 1 and predicts actions that will be taken on the device. Then, an Android agent tries to execute the actions to validate whether the instructions work with the execution information collected. This is an iterative process until the agent either completes the task or is unable to proceed further. Stage 3 is Reranking: representation features of each ¡page, query¿ pair are extracted from execution information and used for reranking the results. In this stage, search results that help yield better execution results are assigned with higher priority in the ranking.
For the simplification purpose, we have utilized pre-trained GPT models in the first two components, which enables us to rapidly establish a research platform that covers all three necessary components. Now we delve into the technical details of each component: Open Domain Instruction Extraction, On-Device Execution, and Reranking.
3.1. Open Domain Instruction Extraction
The task of extracting potentially relevant instructions from candidate web pages is divided into the generation and grounding phases, as shown in stage 1 of Figure 1. This approach enables us to extract instructions from any web page without collecting any training data, i.e. zero-shot learning.
Motivated by recent research using LLM for information extraction (Sun et al., 2011) and (Pomikálek, 2011), we first prompt a LLM to extract relevant instructions from a given web page, otherwise generate ”none”. In the generation phase, for a given search query and each candidate website , we pass in the prompt of the generative model the search query , candidate web page title, and a cleaned HTML document. The cleaned HTML is generated using a parsing algorithm that removes extraneous information, such as javascript and header information, from the web page but still preserves the hierarchical structure of the HTML snippets.
Since generative models tend to paraphrase instructions or even generate their text content unbounded with the original HTML, it is necessary to introduce a grounding mechanism to guarantee the generated instructions match the content in the HTML to minimize hallucinations in the generated content. For the grounding purpose, we first extract all text snippets , from the cleaned HTML with their corresponding XPaths. The first instruction generated from LLM is matched with the closest candidate snippet , using a combination of Faiss semantic (Johnson et al., 2019) similarity and rouge (Lin, 2004) similarity. A typical matching criterion has been illustrated in section 4.3.1. For each subsequent instruction, only snippets having similar XPaths to the previous grounded steps are considered as a matched step candidate. This approach maintains the sequential nature of extracted instructions. If a HTML page has generated instruction(s) grounded to HTML snippets, the relevant instructions can be used for the next execution stage.
3.2. On-Device Execution
The on-device execution agent is designed to directly execute instructions on an end-user device. Simulating the steps on a device provides a more accurate judgement on whether the instructions can be completed on a specific device. In our research, three components are developed to achieve this goal and the workflow is depicted in Figure 2:
Runtime UI Context Detection Module provides UI context information for the current screen state in a JSON format. It contains the instructions, UI control hierarchy, possible actions to take, and metadata for actionable regions. It serves as the input for the Action Prediction Model and is also used to extract reranking features (Sun et al., 2022).
Multimodal Action Prediction Model predicts the action and its required parameters (e.g. UI control index for click action, direction (up, down) for swipe, etc.) to complete an instruction with a chain of actions (Zhang and Zhang, 2023).
Execution Proxy Module takes the predicted action and executes it on device.
The execution loop continues until either all instructions have been finished, or the Action Prediction Model can’t determine the next action, or it reaches the maximum number of steps222In our experiment, the maximum number was set to to avoid an endless execution loop..
3.2.1. Runtime UI Context Detection
This module is designed to gather data about the device’s current User Interface (UI) context. The UI context includes a screenshot of the UI and metadata of ”actionable regions”. An ”actionable region” refers to a region on the screen where actions, such as clicking or scrolling, can be performed. These regions are hierarchically organized and can be associated with a textual description. For example, a region could be a navigation bar, with a back button and title as embedded child regions.
Sometimes, a region doesn’t contain any textual description, but child regions do have more meaningful textural information, which falls into the missing information in terms of action prediction. We use a control distillation process to fill in this missing information, depicted in Figure 3. First, invisible or non-actionable regions are filtered out. Then, descriptions from the nearest non-actionable child regions are consolidated to fill the missing information.
3.2.2. Multi-modal Action Prediction Model
Following Yang et al. (2023b); Zhang and Zhang (2023); Sun et al. (2022), the Action Prediction Model prompts GPT4-V OpenAI et al. (2023); Yang et al. (2023a), where the current UI context and multi-modal action chains (history of UI screenshots paired with each action taken), are passed as input to yield target action from a set of candidate actions333Injecting the UI context into the prompt can help model determine the next action, but if we can use it as an input to train or finetune the model, we expect more accurate action prediction results..
For timestep t , we try to predict the next action to take by prompting GPT4-V. A prompt is constructed using the following parameters and passed into GPT4-V:
: UI context for the current screen.
: List of screenshots of UI screens from previous actions taken.
: Action description for each previous action taken. Each previous action includes both the action name and the chosen actionable region description.
: Set of possible actions based on all kinds of actions currently found in .
: User attributes that aren’t covered in instructions, such as a sample username, password, and location.
3.2.3. Execution Proxy
On-Device Execution Module:
It takes a generated action and its parameters returned from the Action Prediction Model, locates the action to the corresponding region, finally executes the action.
Auxiliary Service Module:
It facilitates the execution information of agents to other modules, helping us analyze and monitor the execution progress.
3.3. Reranking
Execution information that include both textual and visual information (UI screenshot and corresponding actionable regions) is collected as listed in table 1. A feature extraction module converts these raw data into a reranking feature vector. To avoid unnecessary complexity initially, we now use statistical measures of execution information, and left as future work to explore more advanced feature generation techniques, such as time series transformer or feature learning.
Description | |
Search query | Goal description on Android phone for Apps |
Web page | Top 20 Pages retrieved from search engine |
Instructions | Step-by-step instructions from web page |
Mobile screens | UI screen saved per step |
Distilled control list | Distilled control list on screen per step |
Full control list | Full control list on screen per step |
Action & Parameter | Predicted action and its parameter per step |
Action attributing | Which instruction drives model to take action |
Runtime information | More runtime information (issue, start time, etc.) |
3.3.1. Feature Design
Commonly, all textual and visual features are normalized to a float value between 0 and 1. Below is a brief description of our used features.
Id | |
Textual Features | |
Query term ratio in instructions | 1 |
Relevance between page and search query | 2 |
Keyword ratio | 3 |
Visual Features | |
Instructions completion degree | 4 |
Action term ratio in instructions (avg, min, max, var) | 5-8 |
UI term matching ratio in instructions (avg, min, max, var) | 9-12 |
Matched UI term frequency on UI (avg, min, max, var) | 13-16 |
Relative position of the last matched instruction term | 17 |
Moving distancing of instruction terms | 18 |
Textual Features
: Query term frequency ratio in instructions measures the relative frequency of ”How-to” search query terms in the extracted instructions I, refer to Equation 1.
Q is the set of tokens from the search query removed of stopwords, where is the term frequency of token in the extracted instructions I.
: Relevance between page and search query measures the degree of relevance between a web page and the corresponding ”How-to” search query. We prompt GPT to generate a relevance score ranging from [0, 1] using the search query, extracted instructions from the page, and the page title.
: Average term frequency of common keywords in instructions: This feature measures the average frequency of common keywords found in the extracted instructions. We have extracted instructions from about 1000 web pages from Google Android Help Website and identified a set of most common keywords, removing stopwords.
Visual Features
: ”Instructions completion degree” measures the degree to which the extracted instructions have been completed. GPT4-V is prompted to return a completion score ranging from [0, 1], passing the history of UI screenshots for each action taken during execution and extracted instructions as input.
: average, min, max, variance of Action term ratio in instructions measures the average, minimal, maximum, standard deviation of the percentage of instruction words that appear in each action description during the entire execution.
: average, min, max, variance of Visible UI Term Ratio in instructions measures the average, minimal, maximum, standard deviation of the percentage of the current screen UI control’s visible texts appeared in the instructions during the entire execution. This metric indicates the alignment between the current UI screen and the instructions. For example, at the sign-up phase, instructions often reference fields like username, password, and gender. A higher correspondence of these terms with UI elements suggests that the agent is progressing correctly.
: average, min, max, variance of Distilled UI ratio in instructions describes the average, minimal, maximum, standard deviation of the occurrence of distilled UI control text in the instructions during the entire execution. These metrics reflect the likelihood of our execution model accurately selecting the appropriate UI control from all distilled options.
to : These capture how much of each extracted step is executed by comparing the list of actions taken for a given step444A given step from a plan can require multiple actions to complete to its step text.
: Relative Position of Last Matched Instruction Term identifies all matched positions of Action Description in instructions and uses the maximum value to quantify how far our execution can move from the starting position of extracted instructions.
: The moving distancing of instruction terms measures the moving length from the first matched instruction term to the last matched one, indicating the action coverage in instructions.
Due to the page limit, we leave the exact equations for calculating the heuristic features to in our open source web site (TBA).
3.3.2. Reranking Models
In this section, we will illustrate which candidate pages will be extracted out reranking features and the details of our reranking models.
Verified Pages
When the system fails to extract instructions or those pages do not contain any instruction, a candidate page does not progress to the on-device execution stage. Therefore, our reranking module only uses a heuristic to prioritize all verified pages and rerank them subsequently. In our experiments, a page is considered as verified only if it progresses to stage 2: that is, the execution module can successfully execute at least one step and the predicted instruction completion rate is non-zero. All other Pages are kept with their original relative order (i.e. not reranked).
Models
Without losing the generality, three different learning-to-rank(LTR) models can be trained to rerank the verified pages using pre-defined features.
The first model is a simple pairwise LTR logistic regression model, which optimizes the relative order of pairs of pages. Another two types of models are multilayer perception (MLP) transformer models, short for TMLP, driven by two loss functions. One is LambdaLoss (Wang et al., 2018), a pairwise loss function. For each pair of items, the lambda value represents the change in the overall ranking metric, if the order of the two items were swapped. The lambda values are used to weight the errors in the pairwise comparisons, thus this loss function will penalize misorderings more if they have a higher lambda value. Another is NeuralNDCG loss (Qin et al., 2020), a method directly optimizing NDCG metric. This method introduces a NeuralSort operator based on a differentiable function (softmax) to approximate the non-differentiable ground-truth sort operator.
where is the maxDCG at k-th rank, scale(·) is Sinkhorn scaling, g(·) and d(·) are their gain and discount functions.
4. Experiments
Experiments are set up to analyze whether the proposed method can improve the retrieval results of state-of-the-art search engines. We used Google as our original search engine throughout the whole process. We developed MagicWand, a web platform that standardizes the training, execution, and evaluation pipeline for intelligent agents running on Android devices. It can spin up multiple Android emulators and connect with real devices on the server, where each device is controlled by the user on a web interface or ADB-equivalent command interfaces.
4.1. Test Data: ”How-to” WeWeb dataset
A new ”How-to” WeWeb dataset has been collected using MagicWand, which covers instruction extraction, on-device execution, and data annotation for reranking. This dataset is used to measure the end-to-end accuracy of our system.
The finalized dataset version contains queries and instructions for 17 different apps spanning 9 categories of the Google Play Store and an additional ”System” domain for the default system app. We selected a single ”How-to” search query in the System domain. For other domains, we recruit 6 students to propose about ten ”How-to” search queries per app. Using these queries, we retrieved valid web pages from the top 20 Google search results555For the ”Entertainment” domain, there are 18 search results that didn’t have valid web pages, but linked to YouTube videos; For the ”News” category, 1 search result wasn’t a valid web page..
Domain | Application | Queries | Pages | Instructions |
System | Settings | 1 | 20 | 12 |
Entertainment | Pluto TV | 11 | 202 | 9 |
YouTube | 12 | 240 | 26 | |
Education | Coursera | 10 | 200 | 2 |
Quizlet | 10 | 200 | 9 | |
Food & Drinks | DoorDash | 10 | 200 | 4 |
McDonalds | 10 | 200 | 2 | |
Communication | Google Chat | 10 | 200 | 9 |
Messenger | 12 | 240 | 27 | |
Shopping | Target | 10 | 200 | 1 |
eBay | 10 | 200 | 4 | |
News | 10 | 199 | 11 | |
BBC News | 10 | 200 | 1 | |
Maps&Navigation | Google Maps | 10 | 200 | 45 |
Here WeGo | 11 | 220 | 2 | |
Travel | Trip Advisor | 10 | 200 | 4 |
Expedia | 10 | 200 | 0 | |
Total | 167 | 3321 | 168 |
We hired Amazon Mechanical Turk workers to extract relevant instructions from candidate web pages for ”How-to” queries. Multiple workers analyzed each page to ensure data quality, with consensus needed on the extraction results to label as a positive case. The resulting data, detailed in Table 3, includes statistics on the number of search queries, result pages, and pages with valid instructions for each domain.
Human Verification Information:
With search queries and instructions extracted (firstly by humans and further verified by the extraction module) from pages returned by Google, tasks were set up on MagicWand to assess and record the execution of these instructions on Android devices. 16 workers manually read each instruction, executed the instructions on MagicWand, and labeled whether the ”How-to” queries could be completed based on the instructions on each page. Out of all pages evaluated, 168 pages are labeled as successful (i.e. positive) and 64 queries were associated with at least one positively labeled page. These human labels are adopted as our ground truth , which is assigned a value of 1 if a human annotator can follow the instructions on the page and complete the task, otherwise 0.
We captured the actions performed by the workers on the devices, along with their annotations, to the well formatted files. These documents included screenshots, JSON files that detailed the history of actions taken, and comprehensive screen recordings capturing the entire interaction sequence. Additionally, a specialized plugin integrated within the platform was employed to extract the UI context (see Section 3.2.1 having each action executed by the workers).
4.2. Evaluation Metrics
Common reranking metrics for evaluating model performance are used: mean reciprocal rank (MRR) of the top one retrieved documents(MRR@1), precision of the top one and five retrieved documents (P@1 and P@5) and normalized discounted cumulative gain (NDCG) of the top five retrieved documents (NDCG@5).
4.3. System Configurations
4.3.1. Extraction
Generative model:
The state-of-art LLM, GPT4 (OpenAI et al., 2023), has been used to extract instructions from a candidate web page. We set the temperature parameter as 0 to encourage direct extraction and minimize generative content. We also design a prompt containing a list of valid and relevant instruction descriptions. We also try cohere (Documentation, 2024) using the same prompt, which could output accurate citations from the input web page, and found there was no noticeable performance difference compared with GPT4.
Grounding:
A rouge score of above 0.7 or faiss similarity (Johnson et al., 2019) score of under 0.25 is used to match a HTML text snippet of a webpage to a generated step.
4.3.2. Pre-training Reranking Models
Synthetic dataset for pre-training
Since the ”How-to” WeWeb dataset is small, we decided to use the whole dataset for testing; thus the reranking is a zero-shot learning problem. We created synthetic data to pre-train pairwise LTR models. The synthetic pre-training data - ”How-to” META-GUI dataset was created based on META-GUI (Sun et al., 2022), a multimodal dataset for task-oriented conversational agents running on Android. The original data set includes 1125 conversations, 4684 dialogue turns, and 18337 Android execution trajectories. It covers six apps: weather, calendar, search, taxi, restaurant, and hotel mobile apps. This dataset has been used to generate a new ”How-to” Meta-GUI dataset by the following process:
Given 62 few-shot samples(the paired instruction, execution, search query), GPT3.5 is prompted to summarize an execution trajectory with its dialog turn to generate a list of step-by-step instructions and a relevant ”How-to” search query. These serve as positive examples.
Additional positive examples are created by swapping similar search queries in the same app domain from the generated samples of step one.
Negative examples are created by identifying trajectories that fail to complete.
We randomly modify or delete actions in a given trajectory from examples of step one to create further negative samples.
Table 4 summarizes the number of queries, executions, and instructions of our ”How-to” META-GUI dataset.
Query | Executions | Instructions | |
Positive examples | 7884 | 13516 | 4684 |
Negative examples | 944 | 25764 | 4684 |
Total | 8828 | 39280 | 4684 |
This synthetic dataset significantly differs from real data and the testing set. However, we expect possessing some form of pre-training material, even if not ideal, is preferable to having none at all, particularly in the context of zero-shot learning scenarios. We split ”How-to” META-GUI dataset into training, validation, and test sets, then train the LTR models with the most suitable hyperparameters. Next, the trained models are applied to ”How-to” WeWeb dataset.
Reranking model setting
Regarding the model details, TMLP is a fully connected layers of 96-dimensional input without normalization, followed by two transformer blocks with 384-dimensional features and a dropout rate of 0.1. A post-processing model takes the output of the transformer blocks and uses a linear layer to output the reranking score. For NeuralNDCG loss, its temperature is set as 1.0 and we use a ”powered relevancies” gain function , where x is the relevance score. For LambdaRank loss, we set its parameters = 10, and = 1.0. Both are optimized by the same training setting: an Adam optimizer with an initial learning rate of 0.001 is applied with a learning rate scheduler of size 50 and decreasing rate = 0.1. Both training epochs and early stopping patience parameters are set to 20 with NDCG@5 as the validation metric.
4.4. Methods compared
To verify the model performance, we compared the following methods and calculated each query’s MRR, P@1, P@5 and NDCG@5, using as the relevance score.
Oracle: This method consistently ranks the correct pages before the incorrect pages, which gives the theoretical optimal performance.
Baseline Google: We order pages with/without extracted instructions by their Google rank and put the group with relevant instructions ahead of those without.
-based rank: A single feature (i.e. instructions completion degree predicted by GPT4-V) is used to rerank pages with extracted instructions.
LR: This method uses a logistic regression based LTR model, pre-trained on a different synthetic dataset.
NeuralNDCG + TMLP: This method uses a transformer with multilayer perception for LTR, pre-trained using NeuralNDCG loss (Qin et al., 2020) on a different synthetic dataset.
LambdaLoss + TMLP: This method uses a transformer with multilayer perception for LTR, pre-trained using Lambda loss (Wang et al., 2018) on a different synthetic dataset.
5. Experimental Results
5.1. Comparison with Baselines
The performance metrics of different models are shown in Table 5. The ”Oracle” model has the highest performance across all metrics, which is expected as it serves as the ideal benchmark. The baseline (i.e. Google) has the lowest performance, showing that all other models have improved the original ranking. The improvements over the baseline are statistically significant for LR, NeuralNDCG + TMLP, and LambdaLoss + TMLP models. Despite the simple zero-shot learning nature of the models used in our experiments, the results demonstrate adding verification and reranking into the retrieval process can significantly enhance the performance of a leading baseline search engine.
In comparison, -based rank performs worse than other reranking models. This shows reranking using a strong yet single feature is feasible but not good enough. Neural ranking models (TMLP) are better than LR probably because they can capture feature interrelationships. However, the difference is not very big, possibly due to the low quality of the synthetic data used for pre-training.
Results are performance on ”How-to” WeWeb dataset using zero-shot learning. , and are statistically significantly better than the baseline for . Model MRR P@1 P@5 NDCG@5 Oracle 0.3353 0.3353 0.1365 0.3353 Baseline: Google 0.1782 0.1138 0.0850 0.1692 -based rank() 0.2107 0.1737 0.0946 0.1971 LR 0.2406 0.1976 0.1006 0.2227 NeuralNDCG + TMLP 0.2566 0.2275 0.1006 0.2275 LambdaLoss + TMLP 0.2594 0.2335 0.1030 0.2350
5.2. Further Analysis and Ablation Studies
Despite the promising performance shown in the zero-shot setting, we want to further explore the correctness based on our data setting and in which cases the system stumbled. Therefore, we conducted further analysis in this section to answer several critical questions.
Q1. When do our proposed methods work and fail? Why?
To answer this question, an analysis of both success and failure cases are conducted to measure the failure rate at each stage. After the instruction extraction stage, 398 pages (11.98%) have extracted instructions from the total 3321 pages, with 117 queries (about 70.05%) out of the total 167 queries. Among the 398 pages, 75 pages (18.8%) with 39 queries(33.33%) contain relevant instructions, while the remaining 323 pages with 108 queries are negative samples. At this stage, LLMs’ hallucination issue is a major factor: generative content unaligned with original page content induces noises, which decreases the percentage of true positive pages. We will further analyze how the extraction performance has been impacted by HTML grounding in question 2.
At the execution stage, the execution engine filters out 119(29.90%) pages from 68(58.12%) queries and only 279 pages from 91 queries are left. For those remaining pages, if the execution engine can finish more than one instruction or the instruction completion degree of the corresponding execution is larger than 0, the extracted instructions are considered as executable and are forwarded to the next stage. Among those dropped, 18(15.12%) pages are ones with no action taken by the execution engine and 101 (84.88%)pages have a score of 0 generated by GPT4-V. Considering the result of the former is promising, we focus on analysis of : clearly shown in table 6, about 25% of pages with no or partial instructions are set with = 0 and about 70% makes the execution engine struck at the first instruction (the engine can take actions on mobile screen but can’t move forward due to the missing necessary information). Another good indicator is that about 75% pages with instructions can support the execution engine to finish more than 75% instructions, which is aligned with human verification. However, top 25% pages with no or partial instructions are mistakenly marked as = 1, introducing significant noise in reranking. This highlights the necessity of drawing more robust, execution-derived signals from the execution in future work.
Pages | Mean | Std | Min | 25% | 50% | 75% | Max |
With instructions | 0.8647 | 0.2438 | 0 | 0.75 | 1 | 1 | 1 |
No/Partial instructions | 0.3697 | 0.3218 | 0 | 0 | 0.5 | 0.5 | 1 |
At the reranking stage, we analyze the relationship between features and reranking scores for the best-performed models. Most successful cases are because models figure out strong signals among input features, such as and , and learn accurate non-linear weights to reflect on their interrelationship. However, current models tend to minimize negative signals even for those with abnormal values: for instance, a feature vector with several very low values, like 2.8e-05, will still be assigned a higher score just because of multiple features that contain higher values. We attribute this to insufficient negative examples of training and validation sets or limited features incorporated in reranking.
Q2. What’s the performance of the extraction module? Does grounding matter for extraction? How instruction extract works and how it impacts the final results (failed due to extraction failure):
We use page URLs collected in ”How-to” WeWeb dataset to fetch the corresponding web pages and apply instruction extraction on individual pages to compare whether the extracted content matches the ground truth steps. Accuracy is used to evaluate the instruction extraction performance, which measures whether the model appropriately extracts steps or outputs ”none” based on the given web page and search query. GPT4 extraction accuracy is compared with and without step grounding (Table 7). The results show IE without grounding (full or partially) is pages out of 169 web pages with relevant instructions. 42 pages with instructions were partially extracted. However, it also falsely extracted instructions from 953 pages that don’t contain relevant instructions. This happens because the model can hallucinate, extract the wrong set of instructions from a web page with multiple sets of instructions, or significantly paraphrase the instructions.
With grounding, many partially extracted instructions are corrected by adding missing steps on the same XPaths on the HTML DOM tree, and hallucinated instructions are removed. As a result, grounding greatly reduced the number of false positive (i.e. extracted instructions from pages without relevant instructions) from 953 to 303, meanwhile reducing the number of true positive (i.e. extracted instructions from pages with relevant instructions) from to .
Extracted Steps | Has relevant instruction | With grounding | Without grounding |
Full Extraction | True | 78 | 100 |
Partial Extraction | True | 18 | 42 |
Full Extraction | False | 303 | 953 |
No Extraction | True | 73 | 27 |
No Extraction | False | 2849 | 2199 |
Total Pages | 3321 | 3321 |
Q3. Does the zero-shot LTR for reranking help?
We expect LTR will work if we have training data with the same distribution as the test data. However, considering the huge difference between the pre-training and testing data, is it still valid to have a LTR component? We tried to rerank by a simple rule: promoting pages with (i.e. GPT4-V predicts the instructions are 100% completed) to the top without changing their related rankings. This rule has a , significantly worse than the proposed approach and not significantly better than the baseline. This suggests the LTR module is more useful than the simple rule.
Q4. How sensitive the model performance is for different LTR model settings?
We varied the TMLP model structure by adding Sigmoid, Tanh activation functions. We found the model performance changed a little. However, there is nothing statistically significant. We expect the small size of our test dataset and the big gap between training and testing data causes the results less sensitive to the perturbations on TMLP. The LTR models we trained are far from optimal, and we expect better training data could improve the performance.
5.3. Discussion and Future Work
While our study represents a novel approach for search ranking, we recognize that this is a preliminary study of the problem, and as a result comes with certain limitations.
Sample size, scope and experimental constraints:
One of the primary limitations of our study is that it has been done only on the Android platform, and other platforms such as desktop and web browser are not covered and should be studied in the future. Although we cover a diverse set of domains, the coverage is not exhaustive. Our study also shows the exact performance (P@1 etc.) is domain-specific. This may affect the generalization ability of our findings and the reader should take this into account when interpreting the results. When applying similar techniques to other platforms or domains, the exact number (P@1 etc.) would be different.
Extension of instruction extraction:
Although generative LLMs offer a rapid start, the extraction model is significantly hindered by hallucination issues, as noted by Ji et al. (Ji et al., 2023). Our current methodology, which focuses on post-processing with grounding, proves to be a weak regularizer. Moving forward, we would like to try generative LMs with multimodal document models such as XDoc (Chen et al., 2022), MarkupLM (Li et al., 2022b), and LayoutLMv3 (Huang et al., 2022b), aiming to enhance in-model grounding efficacy.
Refinement of execution module:
A more tightly coupled integration of the UI context and the Action Prediction Model can further enhance the execution performance. Moreover, our current prompt-based method which used the actionable regions as input does not effectively represent semantic groups on the UI layer. For instance, an input box and its label username should be semantically grouped as one region for the execution module. Another next step is to enhance the previous work (Hu, 2020), (Liu et al., 2018) to improve the UI contextual information passed to the Action Prediction Model.
Reranking model:
The reranking model we used is very preliminary, using simple feature engineering and no real-world training data. With more resources, we should collect more real-world training data with strategic sampling. Instead of using manually engineered features, we could represent the verification information, especially the sequence of action screens and instructions, through supervised or unsupervised learning. With better reranking models, we would expect the effectiveness of the proposed approach to be more significant.
Safety issues:
Although verification agents can save a user time, it might accidentally take harmful actions as some online content could be misleading. In practice, we need to implement safeguards to prevent harmful actions from being taken by the execution agent. Ideally, a client environment that doesn’t impact the user’s real environment should be used. If not, we might let each user review and approve extracted instructions or at-risk actions before the automated execution. Another approach is to run server-side verification offline and display metadata about verification in the search results (e.g. ”verified for app1 v3 on android v11”), which allows personalized search results without real time verification on the client side.
6. Conclusion
We propose reranking top retrieved results for ”How-to” search queries by promoting candidate web pages with verified executable instructions, which adds an additional layer of usefulness to a typical search engine workflow. It ranks search results based on the actual completeness of the candidate instructions, besides traditional metrics such as textual relevance and authority scores. The experimental results demonstrate the proposed approach could further improve a very strong baseline search engine(i.e. Google).
This paper is a first step forward in improving the reliability and usefulness of search results of online help resources for ”How-to” queries about software tasks. There are several future directions as discussed before, like adapting the system 1 to support verification of other platforms (web, macOS, Windows, iPhone etc.), or collecting user feedback to improve the reranking models and also improving the software agent’s ability to interpret and execute a wider range of instructions.
Acknowledgements.
For ”How-to” WeWeb dataset, we appreciate our research collaborators Sijia Zhong, Valentina Tang, Zoey Zhou, Jimmy Chen, and Chole Wong. Without their efforts, it is impossible to finish collecting execution data in a short time period. They gave valuable suggestions on MagicWand, making it more mature to support our research work. Thank Yaxuan Wang for her efforts to refine figures used in this paper, helping illustrate key ideas concisely and elegantly. Annabelle Miin, Krish Pai, Rithvik Chavali, Saket Pathak have involved this research from the beginning and finished work, like deciding on Android Apps, search queries, and guaranteeing data quality via two phases verification.References
- (1)
- adb library for adb service (2022) Python adb library for adb service. 2022. adbutils. https://github.com/openatx. https://github.com/openatx/adbutils
- Bevendorff et al. (2023) Janek Bevendorff, Sanket Gupta, Johannes Kiesel, and Benno Stein. 2023. An Empirical Comparison of Web Content Extraction Algorithms. (2023).
- Brohan et al. (2022) Anthony Brohan, Noah Brown, Justice Carbajal, Yevgen Chebotar, Joseph Dabis, Chelsea Finn, Keerthana Gopalakrishnan, Karol Hausman, Alex Herzog, Jasmine Hsu, et al. 2022. Rt-1: Robotics transformer for real-world control at scale. arXiv preprint arXiv:2212.06817 (2022).
- Burges et al. (2005) Chris Burges, Tal Shaked, Erin Renshaw, Ari Lazier, Matt Deeds, Nicole Hamilton, and Greg Hullender. 2005. Learning to rank using gradient descent. In Proceedings of the 22nd international conference on Machine learning. 89–96.
- Burges (2010) Christopher JC Burges. 2010. From ranknet to lambdarank to lambdamart: An overview. Learning 11, 23-581 (2010), 81.
- Chen et al. (2022) Jingye Chen, Tengchao Lv, Lei Cui, Cha Zhang, and Furu Wei. 2022. XDoc: Unified Pre-training for Cross-Format Document Understanding. In Findings of the Association for Computational Linguistics: EMNLP 2022. 1006–1016.
- Deka et al. (2017) Biplab Deka, Zifeng Huang, Chad Franzen, Joshua Hibschman, Daniel Afergan, Yang Li, Jeffrey Nichols, and Ranjitha Kumar. 2017. Rico: A mobile app dataset for building data-driven design applications. In Proceedings of the 30th annual ACM symposium on user interface software and technology. 845–854.
- Deng et al. (2022) Xiang Deng, Prashant Shiralkar, Colin Lockard, Binxuan Huang, and Huan Sun. 2022. Dom-lm: Learning generalizable representations for html documents. arXiv preprint arXiv:2201.10608 (2022).
- Documentation (2024) Cohere Documentation. 2024. Cohere Documentation. Cohere AI. https://docs.cohere.com/
- Finn et al. (2001) Aidan Finn, Nicholas Kushmerick, and Barry Smyth. 2001. Fact or Fiction: Content Classification for Digital Libraries.. In DELOS.
- Friedman (2001) Jerome H Friedman. 2001. Greedy function approximation: a gradient boosting machine. Annals of statistics (2001), 1189–1232.
- Grover et al. (2019) Aditya Grover, Eric Wang, Aaron Zweig, and Stefano Ermon. 2019. Stochastic optimization of sorting networks via continuous relaxations. arXiv preprint arXiv:1903.08850 (2019).
- Guideline (2023a) Google Official Developer Guideline. 2023a. Android Developers ¿ Develop ¿ Android Studio ¿ SDK tools Android Debug Bridge (adb). Google. https://developer.android.com/tools/adb
- Guideline (2023b) Google Official Developer Guideline. 2023b. Android Developers ¿ Develop ¿ Reference AccessibilityService. Google. https://developer.android.com/reference/android/accessibilityservice/AccessibilityService
- Guo et al. (2016) Jiafeng Guo, Yixing Fan, Qingyao Ai, and W Bruce Croft. 2016. A deep relevance matching model for ad-hoc retrieval. In Proceedings of the 25th ACM international on conference on information and knowledge management. 55–64.
- Gur et al. (2022) Izzeddin Gur, Ofir Nachum, Yingjie Miao, Mustafa Safdari, Austin Huang, Aakanksha Chowdhery, Sharan Narang, Noah Fiedel, and Aleksandra Faust. 2022. Understanding html with large language models. arXiv preprint arXiv:2210.03945 (2022).
- Hernandez et al. (2021) Alberto Olmo Hernandez, Sarath Sreedharan, and Subbarao Kambhampati. 2021. GPT3-to-plan: Extracting plans from text using GPT-3. CoRR, abs/2106.07131 (2021).
- Hu (2020) Rui Hu. 2020. Detection and Segmentation of Graphical Elements on GUIs for Mobile Apps Based on Deep Learning. In Mobile Computing, Applications, and Services: 11th EAI International Conference, MobiCASE 2020, Shanghai, China, September 12, 2020, Proceedings, Vol. 341. Springer Nature, 187.
- Huang et al. (2022a) Wenlong Huang, Pieter Abbeel, Deepak Pathak, and Igor Mordatch. 2022a. Language models as zero-shot planners: Extracting actionable knowledge for embodied agents. In International Conference on Machine Learning. PMLR, 9118–9147.
- Huang et al. (2022b) Yupan Huang, Tengchao Lv, Lei Cui, Yutong Lu, and Furu Wei. 2022b. LayoutLMv3: Pre-training for Document AI with Unified Text and Image Masking. arXiv preprint arXiv:2204.08387 (2022).
- Ji et al. (2023) Ziwei Ji, Nayeon Lee, Rita Frieske, Tiezheng Yu, Dan Su, Yan Xu, Etsuko Ishii, Ye Jin Bang, Andrea Madotto, and Pascale Fung. 2023. Survey of hallucination in natural language generation. Comput. Surveys 55, 12 (2023), 1–38.
- Johnson et al. (2019) Jeff Johnson, Matthijs Douze, and Hervé Jégou. 2019. Billion-scale similarity search with GPUs. IEEE Transactions on Big Data 7, 3 (2019), 535–547.
- Jung et al. (2021) Geunseong Jung, Sungjae Han, Hansung Kim, Kwanguk Kim, and Jaehyuk Cha. 2021. Don’t read, just look: Main content extraction from web pages using visual features. arXiv preprint arXiv:2110.14164 (2021).
- Lewis et al. (2020) Patrick Lewis, Ethan Perez, Aleksandra Piktus, Fabio Petroni, Vladimir Karpukhin, Naman Goyal, Heinrich Küttler, Mike Lewis, Wen-tau Yih, Tim Rocktäschel, et al. 2020. Retrieval-augmented generation for knowledge-intensive NLP tasks. In Proceedings of the 34th International Conference on Neural Information Processing Systems. 9459–9474.
- Li et al. (2023) Junnan Li, Dongxu Li, Silvio Savarese, and Steven Hoi. 2023. Blip-2: Bootstrapping language-image pre-training with frozen image encoders and large language models. arXiv preprint arXiv:2301.12597 (2023).
- Li et al. (2021) Junlong Li, Yiheng Xu, Lei Cui, and Furu Wei. 2021. Markuplm: Pre-training of text and markup language for visually-rich document understanding. arXiv preprint arXiv:2110.08518 (2021).
- Li et al. (2022b) Junlong Li, Yiheng Xu, Lei Cui, and Furu Wei. 2022b. MarkupLM: Pre-training of Text and Markup Language for Visually Rich Document Understanding. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). 6078–6087.
- Li et al. (2022a) Shuang Li, Xavier Puig, Chris Paxton, Yilun Du, Clinton Wang, Linxi Fan, Tao Chen, De-An Huang, Ekin Akyürek, Anima Anandkumar, et al. 2022a. Pre-trained language models for interactive decision-making. Advances in Neural Information Processing Systems 35 (2022), 31199–31212.
- Li et al. (2020) Yang Li, Jiacong He, Xin Zhou, Yuan Zhang, and Jason Baldridge. 2020. Mapping Natural Language Instructions to Mobile UI Action Sequences. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics. 8198–8210.
- Li et al. (2019) Yuanchun Li, Ziyue Yang, Yao Guo, and Xiangqun Chen. 2019. Humanoid: A deep learning-based approach to automated black-box android app testing. In 2019 34th IEEE/ACM International Conference on Automated Software Engineering (ASE). IEEE, 1070–1073.
- Lin (2004) Chin-Yew Lin. 2004. Rouge: A package for automatic evaluation of summaries. In Text summarization branches out. 74–81.
- Liu et al. (2018) Thomas F Liu, Mark Craft, Jason Situ, Ersin Yumer, Radomir Mech, and Ranjitha Kumar. 2018. Learning design semantics for mobile apps. In Proceedings of the 31st Annual ACM Symposium on User Interface Software and Technology. 569–579.
- Mialon et al. (2023) Grégoire Mialon, Roberto Dessì, Maria Lomeli, Christoforos Nalmpantis, Ram Pasunuru, Roberta Raileanu, Baptiste Rozière, Timo Schick, Jane Dwivedi-Yu, Asli Celikyilmaz, et al. 2023. Augmented language models: a survey. arXiv preprint arXiv:2302.07842 (2023).
- Mitra et al. (2017) Bhaskar Mitra, Fernando Diaz, and Nick Craswell. 2017. Learning to match using local and distributed representations of text for web search. In Proceedings of the 26th international conference on world wide web. 1291–1299.
- OpenAI et al. (2023) OpenAI, :, Josh Achiam, Steven Adler, Sandhini Agarwal, Lama Ahmad, Ilge Akkaya, Florencia Leoni Aleman, Diogo Almeida, Janko Altenschmidt, Sam Altman, and etc. 2023. GPT-4 Technical Report. arXiv:2303.08774 [cs.CL]
- Pobrotyn et al. (2020) Przemyslaw Pobrotyn, Tomasz Bartczak, Mikolaj Synowiec, Radoslaw Bialobrzeski, and Jaroslaw Bojar. 2020. Context-Aware Learning to Rank with Self-Attention. ArXiv abs/2005.10084 (2020).
- Pobrotyn and Bialobrzeski (2021) Przemyslaw Pobrotyn and Radoslaw Bialobrzeski. 2021. NeuralNDCG: Direct Optimisation of a Ranking Metric via Differentiable Relaxation of Sorting. ArXiv abs/2102.07831 (2021).
- Pomikálek (2011) Jan Pomikálek. 2011. Removing boilerplate and duplicate content from web corpora. Disertacnı práce, Masarykova univerzita, Fakulta informatiky (2011).
- Qin and Liu (2013) Tao Qin and Tie-Yan Liu. 2013. Introducing LETOR 4.0 Datasets. CoRR abs/1306.2597 (2013). http://arxiv.org/abs/1306.2597
- Qin et al. (2020) Zhen Qin, Le Yan, Honglei Zhuang, Yi Tay, Rama Kumar Pasumarthi, Xuanhui Wang, Michael Bendersky, and Marc Najork. 2020. Are Neural Rankers still Outperformed by Gradient Boosted Decision Trees?. In International Conference on Learning Representations.
- Radford et al. (2019) Alec Radford, Jeffrey Wu, Rewon Child, David Luan, Dario Amodei, Ilya Sutskever, et al. 2019. Language models are unsupervised multitask learners. OpenAI blog 1, 8 (2019), 9.
- Rawles et al. (2023) Christopher Rawles, Alice Li, Daniel Rodriguez, Oriana Riva, and Timothy Lillicrap. 2023. Android in the wild: A large-scale dataset for android device control. arXiv preprint arXiv:2307.10088 (2023).
- Salehnamadi (2022) Navid Salehnamadi. 2022. Automated Assistive-Service Driven Accessibility Testing for Mobile Applications.
- Salehnamadi et al. (2022) Navid Salehnamadi, Forough Mehralian, and Sam Malek. 2022. Groundhog: An automated accessibility crawler for mobile apps. In 2022 37th IEEE. In ACM International Conference on Automated Software Engineering. IEEE, ACM New York, NY, USA, Rochester, Michigan, USA.
- San et al. (2022) Aidan San, Yuan Zhuang, Jan Bakus, Colin Lockard, David Ciemiewicz, Sandeep Atluri, Yangfeng Ji, Kevin Small, and Heba Elfardy. 2022. PLAtE: A large-scale dataset for list page web extraction. arXiv preprint arXiv:2205.12386 (2022).
- Sun et al. (2011) Fei Sun, Dandan Song, and Lejian Liao. 2011. Dom based content extraction via text density. In Proceedings of the 34th international ACM SIGIR conference on Research and development in Information Retrieval. 245–254.
- Sun et al. (2022) Liangtai Sun, Xingyu Chen, Lu Chen, Tianle Dai, Zichen Zhu, and Kai Yu. 2022. META-GUI: Towards Multi-modal Conversational Agents on Mobile GUI. In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing. 6699–6712.
- Toyama et al. (2021) Daniel Toyama, Philippe Hamel, Anita Gergely, Gheorghe Comanici, Amelia Glaese, Zafarali Ahmed, Tyler Jackson, Shibl Mourad, and Doina Precup. 2021. Androidenv: A reinforcement learning platform for android. arXiv preprint arXiv:2105.13231 (2021).
- Venkatesh et al. (2022) Sagar Gubbi Venkatesh, Partha Talukdar, and Srini Narayanan. 2022. UGIF: UI Grounded Instruction Following. arXiv preprint arXiv:2211.07615 (2022).
- Vogels et al. (2018) Thijs Vogels, Octavian-Eugen Ganea, and Carsten Eickhoff. 2018. Web2text: Deep structured boilerplate removal. In Advances in Information Retrieval: 40th European Conference on IR Research, ECIR 2018, Grenoble, France, March 26-29, 2018, Proceedings 40. Springer, 167–179.
- Wang et al. (2018) Xuanhui Wang, Cheng Li, Nadav Golbandi, Michael Bendersky, and Marc Najork. 2018. The lambdaloss framework for ranking metric optimization. In Proceedings of the 27th ACM international conference on information and knowledge management. 1313–1322.
- Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed Chi, Quoc V Le, Denny Zhou, et al. 2022. Chain-of-thought prompting elicits reasoning in large language models. Advances in Neural Information Processing Systems 35 (2022), 24824–24837.
- Yan et al. (2023) An Yan, Zhengyuan Yang, Wanrong Zhu, Kevin Lin, Linjie Li, Jianfeng Wang, Jianwei Yang, Yiwu Zhong, Julian McAuley, Jianfeng Gao, Zicheng Liu, and Lijuan Wang. 2023. GPT-4V in Wonderland: Large Multimodal Models for Zero-Shot Smartphone GUI Navigation. arXiv:2311.07562 [cs.CV]
- Yang et al. (2023a) Zhengyuan Yang, Linjie Li, Kevin Lin, Jianfeng Wang, Chung-Ching Lin, Zicheng Liu, and Lijuan Wang. 2023a. The Dawn of LMMs: Preliminary Explorations with GPT-4V(ision). arXiv:2309.17421 [cs.CV]
- Yang et al. (2023b) Zhao Yang, Jiaxuan Liu, Yucheng Han, Xin Chen, Zebiao Huang, Bin Fu, and Gang Yu. 2023b. AppAgent: Multimodal Agents as Smartphone Users. arXiv e-prints (2023), arXiv–2312.
- Zhan and Zhang (2023) Zhuosheng Zhan and Aston Zhang. 2023. You only look at screens: Multimodal chain-of-action agents. arXiv preprint arXiv:2309.11436 (2023).
- Zhang and Zhang (2023) Zhuosheng Zhang and Aston Zhang. 2023. You Only Look at Screens: Multimodal Chain-of-Action Agents. arXiv e-prints (2023), arXiv–2309.
- Zitkovich et al. (2023) Brianna Zitkovich, Tianhe Yu, Sichun Xu, Peng Xu, Ted Xiao, Fei Xia, Jialin Wu, Paul Wohlhart, Stefan Welker, Ayzaan Wahid, et al. 2023. Rt-2: Vision-language-action models transfer web knowledge to robotic control. In Conference on Robot Learning. PMLR, 2165–2183.

### v3

@@
-Improving Technical ”How-to” Query Accuracy with Automated Search Results Verification and Reranking
+Enhancing Mobile ”How-to” Queries with Automated Search Results Verification and Reranking
 Abstract.
 Many people use search engines to find online guidance to solve computer or mobile device problems.
-Users frequently encounter challenges in identifying effective solutions from search results, often wasting time trying ineffective solutions that seem relevant yet fail to solve the real problems.
+Users frequently encounter challenges in identifying effective solutions from search results, often wasting time trying ineffective solutions that seem relevant yet fail to solve real problems.
 This paper introduces a novel approach to improving the accuracy and relevance of online technical support search results through automated search results verification and reranking.
-Taking ”How-to” queries specific to on-device execution as a starting point, we first developed a solution that allows an AI agent to interpret and execute step-by-step instructions in the search results in a controlled Android environment.
+Taking ”How-to” queries specific to on-device execution as a starting point, we developed the first solution that allows an AI agent to interpret and execute step-by-step instructions in the search results in a controlled Android environment.
 We further integrated the agent’s findings into a reranking mechanism that orders search results based on the success indicators of the tested solutions.
@@
 1. Introduction
-Often people lack the knowledge to complete a task and rely on a search engine to retrieve potential instructions by asking ”How-to” or similar style queries.
-These requests are ”procedural queries” where the user’s intent is to retrieve step-by-step instructions for performing a specific task or operation.
-However, instructions extracted from retrieved results may not be executable, as the search engine primarily ranks pages based on similiarty of search query and candidate web page (Qin and Liu, 2013).
+When lacking the knowledge to complete a task, people usually rely on a search engine to retrieve potential instructions by asking ”How-to” or similar style queries.
+These requests are ”procedural queries” where the user’s intent is to obtain step-by-step instructions for performing a specific task or operation.
+However, instructions extracted from retrieved results may not be executable, as the search engine primarily ranks pages based on the similiarty between the search query and candidate web pages (Qin and Liu, 2013).
 Furthermore, although user engagement information, such as search logs and page quality indicators like PageRank, are utilized for ranking results in modern search engines, the optimal solution often varies on a user-specific basis.
 The use of different operating systems, screen types, and app versions motivates a personalized reranker.
-In addition, search engines integrated with large language models (LLMs) haven shown to involve the likelihood of producing hallucinations (Lewis et al., 2020).
+In addition, search engines integrated with large language models (LLMs) often unavoidably involve the risk of generating inaccurate or fabricated information, commonly known as hallucinations (Lewis et al., 2020).
 How do users solve the problem?
-Typically, a user could visit multiple retrieved pages.
+Typically, a user could visit multiple retrieved pages from a search engine triggered by a ”How-to” query, which briefly describes their ultimate goal.
 If a page that looks relevant contains some step-by-step instructions, the user might try to follow the instructions to see whether it solves the problem.
-However, users frequently experience frustration and waste considerable time attempting various instructions across different pages.
-Can a search engine assist the user in a more efficient way?
-Motivated by the users’ time-consuming manual solution, we substitute this tedious process with an automatic instruction verification process as an additional component of the search process.
-Our hypothesis is that search results verification for technical ”How-to” queries can achieve decent accuracy given the recent research progress on multimodal LLMs, especially with the help of state-of-the-art GPT agents.
-In particular, we introduce a three-stage solution to do instruction set verification and reranking relevant pages, as illustrated in Figure 1: The first stage involves an Instruction Extraction Model to obtain step-by-step instructions from each retrieved page for a ”How-to” search query.
-The second stage verifies their quality by automatically trying or simulating the instruction execution on devices.
+However, the potential issue is that they often experience frustrations and waste considerable time attempting various instructions across different pages.
+Can a search engine further assist the user more efficiently?
+Inspired by such time-consuming manual procedure, we endeavor to substitute it with an automatic instruction verification process as an additional component of the search process.
+Our hypothesis is that search results verification for technical ”How-to” queries can achieve decent accuracy given the recent research progress on multimodal LLMs, especially with the help of state-of-the-art GPT.
+In particular, we introduce a three-stage solution to verify the instruction set and rerank the relevant pages accordingly, as illustrated in Figure 1: The first stage involves an Instruction Extraction Model to obtain step-by-step instructions from each retrieved page for a ”How-to” search query.
+The second stage verifies their quality by automatically simulating the instruction execution on devices.
 The third stage reranks retrieved results based on the execution information.
-We have developed a comprehensive end-to-end solution to rerank ”How-to” search query results on Android systems and applications across domains.
-This can be enhanced as a platform-agnostic approach with slight adaptions to support other environments, such as web or desktop.
-Our solution enhances search engine results by reranking pages based on a simulated verification on Android devices and preliminary experimental outcomes have demonstrated the effectiveness of our methods.
+As an initiative, we have developed a comprehensive end-to-end solution to rerank ”How-to” search query results on Android systems, selecting mobile applications from different domains.
+This can be enhanced as a platform-agnostic approach with slight adaptions to support other similar environments, such as web, iOS and desktop.
+Our solution enhances search engine results by reranking pages based on simulated verification on Android mobile devices, and our preliminary experimental outcomes have demonstrated the effectiveness of our methods.
 Our key contributions are outlined as follows:
-We propose adding search result verification and reranking into the search process.
+We propose adding search result verification and reranking into the search process, relieving users from tedious manual verification.
 For technical ”How-to” queries, we proposed a three-stage solution.
-At the first stage, we propose an information extraction solution to extract step-by-step instructions from web pages using generative AI with grounding techniques.
-At the second stage, we propose an action agent to take instructions in natural language and execute a series of actions on client device.
-At the third stage, we introduce a range of features that integrate both textual data and the outcomes of executions, and these features are then utilized to rerank the retrieved results.
-To illustrate and support future research in this direction, we developed a new research platform MagicWand 111http://nlp-in-477-l.soe.ucsc.edu:8000/admin/ that can verify and rerank search results for ”How-to” queries across several different domain applications on Android.
-We also collected a new dataset with human annotators.
+In the first stage, we propose an information extraction solution to extract step-by-step instructions from web pages using generative AI with grounding techniques.
+In the second stage, we propose and build out a generic action agent to take instructions in natural language format and execute actions on a client device.
+In the third stage, we introduce a range of features based on the execution outcomes, and further utilize them to rerank the retrieved results.
+To support future research in this direction, we developed a new research platform MagicWand, which facilitates the verification and reranking process for ”How-to” queries of Android mobile applications.
+On top of it, we collected a new ”How-to” WeWeb dataset with human annotators.
 We evaluated the proposed idea and the results show it significantly enhances the performance of a leading baseline search engine (i.e. Google).
@@
 As far as we know, there is no prior work on search results verification for ”How-to” queries using simulated execution.
-On the other hand, the techniques we choose to use when developing different components of the proposed solution are inspired by various existing works from multiple domains.
+Furthermore, the techniques we choose to use when developing different components of the proposed solution are motivated by existing works from multiple domains.
 Therefore, we briefly review three important domains that influenced the technical choices we made in our research.
 2.1. Information Extraction
-Information extraction from web pages (Bevendorff et al., 2023), as a broader concept of our extraction task, was studied heavily back in 1990s (Finn et al., 2001).
+Information extraction from web pages (Bevendorff et al., 2023), as a broader concept of our extraction task, was studied extensively back in the 1990s (Finn et al., 2001).
 There are two methodologies: heuristics based approaches and machine learning based approaches.
@@
 DOM-LM (Deng et al., 2022), MarkupLM (Li et al., 2021) used transformer-based architecture to represent the DOM tree and further improve classification accuracy.
-A new dataset PLAtE of list-like websites, such as online stores, was developed by Amazon research team to analyze how LLMs work on structure extraction (San et al., 2022).
+A new dataset PLAtE of list-like websites, such as online stores, was developed by Amazon research team to analyze how LLMs work in structure extraction (San et al., 2022).
 Researchers also endeavored to better utilize pre-trained LLMs to extract target information to avoid training LLMs.
@@
 Enabling intelligent agents to behave according to instructions involves two main tasks: the first is empowering agents with sufficient knowledge to make decisions and act skillfully.
-The second is executing actions, collecting feedback, and recognizing environmental changes caused by these actions.
+The second is to execute actions, collect feedback, and recognize environmental changes caused by these actions.
 To achieve the first, researchers often leverage and enhance LLMs’ reasoning and tooling capabilities (Mialon et al., 2023).
@@
 Recent studies include using pre-trained GPT2 (Radford et al., 2019) in household simulations (Li et al., 2022a), RT1 (Brohan et al., 2022) and RT2 (Zitkovich et al., 2023) in robot control.
-This section will focus on how to adapt these tasks to the Android platform, on which we will carry out our experiments.
+This section will focus on how to adapt these tasks to the Android platform on which we will carry out our experiments.
 Acting Agent on Android:
@@
 Behavior Cloning and Simple action Chain of Thoughts (Wei et al., 2022)) on the AWIT dataset (Rawles et al., 2023).
-However, they overlooked the Android UI Control hierarchy, and the transformers trained without such knowledge can not predict actions accurately given ambiguous or sparse instructions.
+However, they overlooked the Android UI Control hierarchy, and transformers trained without such knowledge cannot accurately predict actions given ambiguous or sparse instructions.
 (Huang et al., 2022a) demonstrated that LLMs equipped with domain knowledge could break down high-level tasks into mid-level plans.
-Researchers were also starting to explore multi-modal LLM agents in Android automation, as evidenced by Yan et al.
+Researchers were also starting to explore multimodal LLM agents in Android automation, as evidenced by Yan et al.
 (2023); Yang et al.
 (2023b).
-Our Action Prediction Model aligns with this line of research, however, we’ve refined it by distilling the UI Control list from runtime metadata and prompting GPT4-V based on the context.
+Our Action Prediction Model aligns with this line of research, however, we’ve refined them in a more fine-grained manner by distilling the UI Control list from runtime metadata and prompting GPT4-V dynamically based on the context.
 Action Execution:
@@
 (2022); Salehnamadi (2022).
-ADB-based solutions, like AndroidEnv (Toyama et al., 2021), adbutils (adb library for adb service, 2022), need additional connection settings for Android devices and the paired computer, which is unrealistic for end users.
-In comparison, Accessibility Service-based solutions can be delivered as standalone mobile applications due to the availability of Accessibility Service API on most systems.
-In our experiments, we used the Accessibility Service API.
+ADB-based solutions, like AndroidEnv (Toyama et al., 2021), adbutils (adb library for adb service, 2022), need additional connection settings for Android devices and the paired computer, which is unrealistic for end users during daily usage.
+In comparison, Accessibility Service-based solutions can be delivered as standalone mobile applications due to the general availability in most systems.
+Considering this benefit, we decided to utilize the Accessibility Service API to develop our execution proxy, which laid down a solid foundation for action execution.
 2.3. Reranking
 In information retrieval (IR), reranking is the adjustment of document rankings based on query-document relevance after the initial retrieval.
-The evolution of reranking began with RankNet (Burges et al., 2005) in 2005, applying gradient descent in the process of learning a ranking function.
+An important evolution of reranking began with RankNet (Burges et al., 2005) in 2005, applying gradient descent in the process of learning a ranking function.
 LambdaRank (Burges, 2010) later refined this approach by optimizing the gradients of the model scores and addressing issues caused by discontinuous gradients.
@@
 Learning-to-Rank Challenge.
-Subsequently, neural ranking models like DRMM (Guo et al., 2016) and Duet (Mitra et al., 2017) emerged.
+Subsequently, similar neural ranking models like DRMM (Guo et al., 2016) and Duet (Mitra et al., 2017) emerged.
 Recently, (Pobrotyn et al., 2020) introduced a context-aware neural network for item relevance calculation, and (Qin et al., 2020) proposed DASALC, a neural LTR model yielding performance comparable to LambdaMART.
@@
 3. Methods
-The flowchart in Figure 1 outlines our proposed process to improve the accuracy and usefulness of ”How-to” search query results.
-It achieves this by extracting instructions from web pages and validating them through execution on the device.
-The validation outcomes are then used to refine and adjust the search result rankings.
-The system can overcome some limitations of traditional ranking algorithms by incorporating user-specific information (device, OS version, application version, etc.) and collecting verification results from client devices.
+The flow chart in Figure 1 outlines the overall process of our proposal primarily by extracting instructions from Web pages and validating them through the execution on the device.
+The validation outcomes are then used to refine the search result rankings.
+The system can overcome some limitations of traditional ranking algorithms by incorporating user-specific information (device, OS version, application version, etc.) and collecting client-side verification results.
 This is a three-stage process.
-Stage 1 is Instruction Extraction: given the initial top N retrieved results, a generative AI model processes each document to generate a sequence of instructions.
+Stage 1 is Instruction Extraction: we take a high-level goal for an Android application as a ”How-to” search query to retrieve results from a search engine.
+Given the initial top N retrieved web pages, a generative AI model processes each document to generate a sequence of instructions.
 Then, a grounding module analyzes the structure of the HTML web pages and further improves the quality of the extracted instructions.
-Stage 2 is On-device Execution: the Action Prediction Model reads the instructions extracted from stage 1 and predicts actions that will be taken on the device.
-Then, an Android agent tries to execute the actions to validate whether the instructions work with the execution information collected.
-This is an iterative process until the agent either completes the task or is unable to proceed further.
-Stage 3 is Reranking: representation features of each ¡page, query¿ pair are extracted from execution information and used for reranking the results.
-In this stage, search results that help yield better execution results are assigned with higher priority in the ranking.
-For the simplification purpose, we have utilized pre-trained GPT models in the first two components, which enables us to rapidly establish a research platform that covers all three necessary components.
+Stage 2 is On-device Execution: the Action Prediction Model reads the instructions extracted from stage 1 and predicts actions to be taken on the device.
+Next, an Android agent takes those actions to validate whether the instructions work, collecting the execution information simultaneously.
+This is an iterative process until the agent either completes the task or can’t proceed further.
+Stage 3 is Reranking: representation features of each ¡page, query¿ pair are extracted from execution information and used for reranking pages for each query.
+In this stage, search results that help yield better execution results are assigned with higher priority during the ranking.
+For simplification, we have utilized pre-trained GPT models in our implementations, which enables us to rapidly set up a pipeline that covers all three necessary components.
 Now we delve into the technical details of each component: Open Domain Instruction Extraction, On-Device Execution, and Reranking.
 3.1. Open Domain Instruction Extraction
-The task of extracting potentially relevant instructions from candidate web pages is divided into the generation and grounding phases, as shown in stage 1 of Figure 1.
-This approach enables us to extract instructions from any web page without collecting any training data, i.e.
-zero-shot learning.
-Motivated by recent research using LLM for information extraction (Sun et al., 2011) and (Pomikálek, 2011), we first prompt a LLM to extract relevant instructions from a given web page, otherwise generate ”none”.
+This stage has been divided into the generation and grounding phases, as shown in stage 1 of figure 1.
+Our approach allows us to extract instructions from any web page without collecting any training data, i.e.
+zero-shot learning 111By realizing the non-trivial effects caused by the hallucination issue, we have further refined the aforementioned strategy, making the side effect to the minimal level.
+More details will be covered in the discussion section, as well as our future work.
+Following Sun et al.
+(2011); Pomikálek (2011) using LLM for information extraction, we first prompt a LLM model to extract relevant instructions from a given web page, otherwise generate ”none”.
 In the generation phase, for a given search query and each candidate website , we pass in the prompt of the generative model the search query , candidate web page title, and a cleaned HTML document.
 The cleaned HTML is generated using a parsing algorithm that removes extraneous information, such as javascript and header information, from the web page but still preserves the hierarchical structure of the HTML snippets.
-Since generative models tend to paraphrase instructions or even generate their text content unbounded with the original HTML, it is necessary to introduce a grounding mechanism to guarantee the generated instructions match the content in the HTML to minimize hallucinations in the generated content.
-For the grounding purpose, we first extract all text snippets , from the cleaned HTML with their corresponding XPaths.
+Since generative models tend to paraphrase instructions or even generate text content unbounded with the original HTML, we also introduce a grounding mechanism to guarantee the generated instructions match the content in the HTML to minimize hallucinations in the generated content.
+For grounding, we first extract all text snippets , from the cleaned HTML with their corresponding XPaths.
 The first instruction generated from LLM is matched with the closest candidate snippet , using a combination of Faiss semantic (Johnson et al., 2019) similarity and rouge (Lin, 2004) similarity.
 A typical matching criterion has been illustrated in section 4.3.1.
-For each subsequent instruction, only snippets having similar XPaths to the previous grounded steps are considered as a matched step candidate.
-This approach maintains the sequential nature of extracted instructions.
-If a HTML page has generated instruction(s) grounded to HTML snippets, the relevant instructions can be used for the next execution stage.
+For each subsequent instruction, only snippets having similar XPaths to the previous grounded steps are considered as matched step candidates.
+This approach maintains the sequential order of extracted instructions from the original web page.
+If generated instruction(s) can be grounded to HTML snippets from the original web page, the relevant instructions can be used for the next execution stage.
 3.2. On-Device Execution
 The on-device execution agent is designed to directly execute instructions on an end-user device.
-Simulating the steps on a device provides a more accurate judgement on whether the instructions can be completed on a specific device.
-In our research, three components are developed to achieve this goal and the workflow is depicted in Figure 2:
-Runtime UI Context Detection Module provides UI context information for the current screen state in a JSON format.
-It contains the instructions, UI control hierarchy, possible actions to take, and metadata for actionable regions.
-It serves as the input for the Action Prediction Model and is also used to extract reranking features (Sun et al., 2022).
+Simulating the steps on a device provides more accurate judgements on whether the instructions can be completed using a specific device.
+In our research, three components are developed to achieve this goal with the workflow depicted in figure 2:
+Runtime UI Context Detection Module delivers UI context information for the current screen state in a JSON format.
+It contains instructions, the UI control hierarchy, potential actions, and necessary runtime metadata.
+This information serves as input for the Action Prediction Model and is also used to extract reranking features (Sun et al., 2022).
 Multimodal Action Prediction Model predicts the action and its required parameters (e.g.
 UI control index for click action, direction (up, down) for swipe, etc.) to complete an instruction with a chain of actions (Zhang and Zhang, 2023).
-Execution Proxy Module takes the predicted action and executes it on device.
+Execution Proxy Module takes the predicted action and executes it on the client device.
 The execution loop continues until either all instructions have been finished, or the Action Prediction Model can’t determine the next action, or it reaches the maximum number of steps222In our experiment, the maximum number was set to to avoid an endless execution loop..
 3.2.1. Runtime UI Context Detection
-This module is designed to gather data about the device’s current User Interface (UI) context.
-The UI context includes a screenshot of the UI and metadata of ”actionable regions”.
+This module collects data that characterizes the device’s current User Interface (UI) context.
+The UI context consists of a UI screenshot and metadata of ”actionable regions”.
 An ”actionable region” refers to a region on the screen where actions, such as clicking or scrolling, can be performed.
-These regions are hierarchically organized and can be associated with a textual description.
+These regions are hierarchically organized and can be associated with several control properties, like textual description, coordinate, scrollable, etc.
 For example, a region could be a navigation bar, with a back button and title as embedded child regions.
-Sometimes, a region doesn’t contain any textual description, but child regions do have more meaningful textural information, which falls into the missing information in terms of action prediction.
-We use a control distillation process to fill in this missing information, depicted in Figure 3.
+To simplify, we currently focus on leveraging the textual description of the control directly associated with an ”actionable region”, like title, text description, etc.
+Sometimes, a region may not directly contain any textual description, but its child regions often possess meaningful textural information, which is crucial for accurate action prediction but is unfortunately missing.
+We designate a control distillation algorithm to fill in this missing information, depicted in figure 3.
 First, invisible or non-actionable regions are filtered out.
-Then, descriptions from the nearest non-actionable child regions are consolidated to fill the missing information.
-3.2.2. Multi-modal Action Prediction Model
+Then, descriptions from the nearest non-actionable child regions are consolidated hierarchically to represent the missing information.
+3.2.2. Multimodal Action Prediction Model
 Following Yang et al.
 (2023b); Zhang and Zhang (2023); Sun et al.
-(2022), the Action Prediction Model prompts GPT4-V OpenAI et al.
+(2022), the Action Prediction Model prompts a multimodal LLM, GPT4-V OpenAI et al.
 (2023); Yang et al.
-(2023a), where the current UI context and multi-modal action chains (history of UI screenshots paired with each action taken), are passed as input to yield target action from a set of candidate actions333Injecting the UI context into the prompt can help model determine the next action, but if we can use it as an input to train or finetune the model, we expect more accurate action prediction results..
-For timestep t , we try to predict the next action to take by prompting GPT4-V. A prompt is constructed using the following parameters and passed into GPT4-V:
+(2023a) in our current setting, where the current UI context and multimodal action chains 333That is history of UI screenshots paired with each action taken are passed as input to choose the target action from a set of candidate actions 444Injecting the UI context into the prompt can help model better determine the next action based on our primary experiment.
+We expect more accurate action prediction results, if we can use this information appropriately as an input to train or finetune the model., such as click, swipe, input, back, etc.
+For timestep t , GPT4-V will predict the next action given a prompt constructed using the following parameters:
 : UI context for the current screen.
-: List of screenshots of UI screens from previous actions taken.
-: Action description for each previous action taken. Each previous action includes both the action name and the chosen actionable region description.
-: Set of possible actions based on all kinds of actions currently found in .
-: User attributes that aren’t covered in instructions, such as a sample username, password, and location.
+: List of screenshots of UI screens from previous actions.
+: Action description for each previous action, including the action name and the chosen actionable region description.
+: Set of possible actions based on available properties of distilled controls in .
+: Information not covered by instructions but required for agent execution.
+For instance, a specific username and password are needed, when the instruction is ”enter user name and password when open WeChat”.
 3.2.3. Execution Proxy
 On-Device Execution Module:
-It takes a generated action and its parameters returned from the Action Prediction Model, locates the action to the corresponding region, finally executes the action.
+It takes a predicted action and its parameters returned from the Action Prediction Model, locates the action to the corresponding region, and finally executes it.
+Considering a mobile application context can change dynamically, we introduce a heuristic to validate the predicted control associated with the target action and elegantly find a suitable alternative when mismatched.
+For flexibility, We also designate a fallback mechanism for candidate actions, so some of them can adjust slightly given different application settings.
+For instance, a click action can be tried with a similar gesture under instant Message application settings.
 Auxiliary Service Module:
-It facilitates the execution information of agents to other modules, helping us analyze and monitor the execution progress.
+It synchronizes the execution information of the mobile agents to external modules, helping to analyze and monitor the execution progress.
+In addition, this module also implements analysis plugins to integrate with GPT4V and GPT4 to summarize and evaluate the status of each execution trajectory.
 3.3. Reranking
-Execution information that include both textual and visual information (UI screenshot and corresponding actionable regions) is collected as listed in table 1.
-A feature extraction module converts these raw data into a reranking feature vector.
-To avoid unnecessary complexity initially, we now use statistical measures of execution information, and left as future work to explore more advanced feature generation techniques, such as time series transformer or feature learning.
+Execution information, both textual and visual information (UI screenshot and corresponding actionable regions), has been collected as listed in table 1 and converted into a reranking feature vector.
+To avoid unnecessary complexity initially, we now use statistical measures of execution information, and leave it as future work to explore more advanced feature generation techniques, such as time series transformer or feature learning.
 | Description | |
@@
 3.3.1. Feature Design
-Commonly, all textual and visual features are normalized to a float value between 0 and 1. Below is a brief description of our used features.
+All textual and visual features are normalized to a float value between 0 and 1. Below is a brief description of them.
 | Id | |
@@
 : Query term frequency ratio in instructions measures the relative frequency of ”How-to” search query terms in the extracted instructions I, refer to Equation 1.
-Q is the set of tokens from the search query removed of stopwords, where is the term frequency of token in the extracted instructions I.
+Q is the set of tokens from the search query without stopword, where is the term frequency of token in the extracted instructions I.
 : Relevance between page and search query measures the degree of relevance between a web page and the corresponding ”How-to” search query.
@@
 : Average term frequency of common keywords in instructions: This feature measures the average frequency of common keywords found in the extracted instructions.
-We have extracted instructions from about 1000 web pages from Google Android Help Website and identified a set of most common keywords, removing stopwords.
+We have used instructions from about 1000 web pages from Google Android Help Website and identified a set of most common keywords, excluding stopwords.
 Visual Features
 : ”Instructions completion degree” measures the degree to which the extracted instructions have been completed.
-GPT4-V is prompted to return a completion score ranging from [0, 1], passing the history of UI screenshots for each action taken during execution and extracted instructions as input.
-: average, min, max, variance of Action term ratio in instructions measures the average, minimal, maximum, standard deviation of the percentage of instruction words that appear in each action description during the entire execution.
+GPT4-V is prompted to return a completion score of [0, 1], using the history of UI screenshots for each action taken during execution and extracted instructions as input.
+: average, min, max, variance of Action term ratio in instructions measures the average, minimal, maximum, standard deviation of the percentage of instruction’s keyword that appears in each action description during the entire execution.
 : average, min, max, variance of Visible UI Term Ratio in instructions measures the average, minimal, maximum, standard deviation of the percentage of the current screen UI control’s visible texts appeared in the instructions during the entire execution.
 This metric indicates the alignment between the current UI screen and the instructions.
-For example, at the sign-up phase, instructions often reference fields like username, password, and gender.
-A higher correspondence of these terms with UI elements suggests that the agent is progressing correctly.
-: average, min, max, variance of Distilled UI ratio in instructions describes the average, minimal, maximum, standard deviation of the occurrence of distilled UI control text in the instructions during the entire execution.
+For example, for user registration, instructions often contain reference fields like username, password, and gender.
+A higher correspondence of these terms with UI elements suggests that the agent is more likely to move forward along the instructions.
+: average, min, max, variance of Distilled UI ratio in instructions describes the average, minimal, maximum, standard deviation of the occurrence of distilled UI control text in instructions during the entire execution.
 These metrics reflect the likelihood of our execution model accurately selecting the appropriate UI control from all distilled options.
-to : These capture how much of each extracted step is executed by comparing the list of actions taken for a given step444A given step from a plan can require multiple actions to complete to its step text.
+to : These capture how much of each extracted step is executed by comparing the list of actions taken for a given step555One step in a plan may require multiple actions to complete to its step text.
 : Relative Position of Last Matched Instruction Term identifies all matched positions of Action Description in instructions and uses the maximum value to quantify how far our execution can move from the starting position of extracted instructions.
 : The moving distancing of instruction terms measures the moving length from the first matched instruction term to the last matched one, indicating the action coverage in instructions.
-Due to the page limit, we leave the exact equations for calculating the heuristic features to in our open source web site (TBA).
+Due to the page limit, we leave the exact equations for calculating the heuristic features to in our open-source website (TBA).
 3.3.2. Reranking Models
-In this section, we will illustrate which candidate pages will be extracted out reranking features and the details of our reranking models.
+In this section, we will illustrate which candidate pages have to be processed for reranking features and the details of reranking models.
 Verified Pages
-When the system fails to extract instructions or those pages do not contain any instruction, a candidate page does not progress to the on-device execution stage.
-Therefore, our reranking module only uses a heuristic to prioritize all verified pages and rerank them subsequently.
-In our experiments, a page is considered as verified only if it progresses to stage 2: that is, the execution module can successfully execute at least one step and the predicted instruction completion rate is non-zero.
-All other Pages are kept with their original relative order (i.e.
-not reranked).
+When the system fails to extract instructions or those pages do not contain any instruction, a candidate page does not advance to the on-device execution stage.
+After that, our reranking modules prioritize all verified pages using a heuristic.
+In our experiments, a page is considered as verified only if it progresses to stage 2 and satisfies the following criteria: the execution proxy must successfully execute at least one step and the predicted instruction completion rate is non-zero.
+All other Pages are kept in their original relative order (i.e., not reranked) and put after the verified pages.
 Models
-Without losing the generality, three different learning-to-rank(LTR) models can be trained to rerank the verified pages using pre-defined features.
-The first model is a simple pairwise LTR logistic regression model, which optimizes the relative order of pairs of pages.
-Another two types of models are multilayer perception (MLP) transformer models, short for TMLP, driven by two loss functions.
+Without losing the generality, three different learning-to-rank(LTR) models have been trained to rerank the verified pages using the aforementioned features.
+The first model is a simple pairwise LTR logistic regression model, which optimizes the relative order of each pair of pages.
+Another two types of models are multilayer perception (MLP) transformer models, short for TMLP, using two different loss functions.
 One is LambdaLoss (Wang et al., 2018), a pairwise loss function.
@@
 4. Experiments
-Experiments are set up to analyze whether the proposed method can improve the retrieval results of state-of-the-art search engines.
+Experiments have been designed to assess whether the proposed method enhances the retrieval performance of state-of-the-art search engines.
 We used Google as our original search engine throughout the whole process.
-We developed MagicWand, a web platform that standardizes the training, execution, and evaluation pipeline for intelligent agents running on Android devices.
-It can spin up multiple Android emulators and connect with real devices on the server, where each device is controlled by the user on a web interface or ADB-equivalent command interfaces.
+Our experiments are hosted on MagicWand, a web platform that standardizes the training, execution, and evaluation pipeline for intelligent agents running on Android devices.
+It can spin up multiple Android emulators or connect with real devices on the server, which can be controlled by the user on a web interface or ADB-equivalent command interface.
 4.1. Test Data: ”How-to” WeWeb dataset
 A new ”How-to” WeWeb dataset has been collected using MagicWand, which covers instruction extraction, on-device execution, and data annotation for reranking.
-This dataset is used to measure the end-to-end accuracy of our system.
-The finalized dataset version contains queries and instructions for 17 different apps spanning 9 categories of the Google Play Store and an additional ”System” domain for the default system app.
+This dataset is used to measure the end-to-end reranking accuracy.
+Its finalized version contains queries and instructions for 17 different apps spanning 9 categories of the Google Play Store and an additional ”System” domain for the default system app 666Based on Google result, the default system app category has higher proportion of web pages with instructions in top 20 result.
+In comparison, other categories vary for instruction proportion and therefore deserve more attention.
 We selected a single ”How-to” search query in the System domain.
 For other domains, we recruit 6 students to propose about ten ”How-to” search queries per app.
-Using these queries, we retrieved valid web pages from the top 20 Google search results555For the ”Entertainment” domain, there are 18 search results that didn’t have valid web pages, but linked to YouTube videos; For the ”News” category, 1 search result wasn’t a valid web page..
+Using these queries, we retrieved valid web pages from the top 20 Google search results777For the ”Entertainment” domain, there are 18 search results that didn’t have any valid web page, but linked to YouTube videos; For the ”News” category, 1 search result wasn’t a valid web page..
 | Domain | Application | Queries | Pages | Instructions |
@@
 We hired Amazon Mechanical Turk workers to extract relevant instructions from candidate web pages for ”How-to” queries.
-Multiple workers analyzed each page to ensure data quality, with consensus needed on the extraction results to label as a positive case.
+Multiple workers analyzed each page to ensure data quality, requiring consensus on the extraction results to be labeled as positive.
 The resulting data, detailed in Table 3, includes statistics on the number of search queries, result pages, and pages with valid instructions for each domain.
 Human Verification Information:
-With search queries and instructions extracted (firstly by humans and further verified by the extraction module) from pages returned by Google, tasks were set up on MagicWand to assess and record the execution of these instructions on Android devices.
-16 workers manually read each instruction, executed the instructions on MagicWand, and labeled whether the ”How-to” queries could be completed based on the instructions on each page.
+With search queries and instructions extracted (firstly by humans and further verified by the extraction module) from pages, tasks were set up on MagicWand to record and access the execution of instructions on Android devices.
+16 workers read each instruction, executed the instructions manually, and labeled whether the assigned ”How-to” queries could be completed.
 Out of all pages evaluated, 168 pages are labeled as successful (i.e.
 positive) and 64 queries were associated with at least one positively labeled page.
-These human labels are adopted as our ground truth , which is assigned a value of 1 if a human annotator can follow the instructions on the page and complete the task, otherwise 0.
-We captured the actions performed by the workers on the devices, along with their annotations, to the well formatted files.
-These documents included screenshots, JSON files that detailed the history of actions taken, and comprehensive screen recordings capturing the entire interaction sequence.
-Additionally, a specialized plugin integrated within the platform was employed to extract the UI context (see Section 3.2.1 having each action executed by the workers).
+These human labels are used as the ground truth : it is assigned with a value of 1 if a human annotator can follow the instructions on the page and complete the task, otherwise 0.
+We captured the actions performed by the workers, along with their annotations, and saved them into documents with a pre-defined format.
+These documents included screenshots, JSON files that detailed the history of actions taken, and screen recordings capturing the entire interaction sequence.
+Additionally, a specialized plugin integrated with the platform was employed to extract the UI context.
 4.2. Evaluation Metrics
-Common reranking metrics for evaluating model performance are used: mean reciprocal rank (MRR) of the top one retrieved documents(MRR@1), precision of the top one and five retrieved documents (P@1 and P@5) and normalized discounted cumulative gain (NDCG) of the top five retrieved documents (NDCG@5).
+Common reranking metrics are used: mean reciprocal rank (MRR) of the top one retrieved documents(MRR@1), precision of the top one and five retrieved documents (P@1 and P@5) and normalized discounted cumulative gain (NDCG) of the top five retrieved documents (NDCG@5).
 4.3. System Configurations
@@
 Generative model:
-The state-of-art LLM, GPT4 (OpenAI et al., 2023), has been used to extract instructions from a candidate web page.
-We set the temperature parameter as 0 to encourage direct extraction and minimize generative content.
+The state-of-the-art LLM, GPT4 (OpenAI et al., 2023), is used to extract instructions from a candidate web page in our experiment, following Sarkhel et al.
+(2023); Sood and Dandapat (2023).
+We set the temperature parameter as 0 to encourage direct extraction with minimizing generative content.
 We also design a prompt containing a list of valid and relevant instruction descriptions.
-We also try cohere (Documentation, 2024) using the same prompt, which could output accurate citations from the input web page, and found there was no noticeable performance difference compared with GPT4.
+For comparison, we have also tested cohere (Documentation, 2024) using the same prompt, which could output accurate citations from the input web page, and found no noticeable performance difference compared with GPT4.
 Grounding:
 A rouge score of above 0.7 or faiss similarity (Johnson et al., 2019) score of under 0.25 is used to match a HTML text snippet of a webpage to a generated step.
+4.3.2. On-Device Execution
+Device setup
+On MagicWand, we have reserved an Android emulator to run tasks given instructions and collect execution statistics 888Realistically, multiple emulators with various settings better simulate realistic cases.
+However, one device can help simplify our initial work and device maintenance..
+For efficiency, a batch-supported execution engine based on adbutils (adb library for adb service, 2022) is also prepared to run multiple tasks on different devices simultaneously.
+Multimodal action prediction model
+To avoid invalid action parameters and keep exploring capacity, GPT4-V parameter ”temperature” is set to 0.3 together with ”max_tokens” of 300.
+We specify instructions, historical actions and UI screenshots in the prompt and adjust the prompt according to distilled UI controls.
+Execution proxy
+The UI context representation data with the execution status are saved into the central storage with a JSON index file describing the execution and associated resources.
+A GPT4V python plugin was developed to evaluate the execution information and generate the task completion rate for each execution.
 4.3.3. Pre-training Reranking Models
 Synthetic dataset for pre-training
-Since the ”How-to” WeWeb dataset is small, we decided to use the whole dataset for testing; thus the reranking is a zero-shot learning problem.
-We created synthetic data to pre-train pairwise LTR models.
+As the ”How-to” WeWeb dataset is small, we decided to use it for testing; thus, the reranking is a zero-shot learning problem.
+We created another synthetic data to pre-train pairwise LTR models.
 The synthetic pre-training data - ”How-to” META-GUI dataset was created based on META-GUI (Sun et al., 2022), a multimodal dataset for task-oriented conversational agents running on Android.
@@
 Given 62 few-shot samples(the paired instruction, execution, search query), GPT3.5 is prompted to summarize an execution trajectory with its dialog turn to generate a list of step-by-step instructions and a relevant ”How-to” search query.
-These serve as positive examples.
+These are positive examples.
 Additional positive examples are created by swapping similar search queries in the same app domain from the generated samples of step one.
@@
 This synthetic dataset significantly differs from real data and the testing set.
-However, we expect possessing some form of pre-training material, even if not ideal, is preferable to having none at all, particularly in the context of zero-shot learning scenarios.
+However, we expect possessing some form of pre-training material, even if not ideal, is preferable to having none at all, particularly in the zero-shot learning settings.
 We split ”How-to” META-GUI dataset into training, validation, and test sets, then train the LTR models with the most suitable hyperparameters.
@@
 Reranking model setting
-Regarding the model details, TMLP is a fully connected layers of 96-dimensional input without normalization, followed by two transformer blocks with 384-dimensional features and a dropout rate of 0.1.
+Regarding the model details, TMLP is a fully connected layer of 96-dimensional input without normalization, followed by two transformer blocks with 384-dimensional features and a dropout rate of 0.1.
 A post-processing model takes the output of the transformer blocks and uses a linear layer to output the reranking score.
@@
 For LambdaRank loss, we set its parameters = 10, and = 1.0.
-Both are optimized by the same training setting: an Adam optimizer with an initial learning rate of 0.001 is applied with a learning rate scheduler of size 50 and decreasing rate = 0.1.
+They are all optimized in the same training setting: an Adam optimizer with an initial learning rate of 0.001 is applied with a learning rate scheduler of size 50 and a decreasing rate = 0.1.
 Both training epochs and early stopping patience parameters are set to 20 with NDCG@5 as the validation metric.
@@
 -based rank: A single feature (i.e. instructions completion degree predicted by GPT4-V) is used to rerank pages with extracted instructions.
-LR: This method uses a logistic regression based LTR model, pre-trained on a different synthetic dataset.
-NeuralNDCG + TMLP: This method uses a transformer with multilayer perception for LTR, pre-trained using NeuralNDCG loss (Qin et al., 2020) on a different synthetic dataset.
-LambdaLoss + TMLP: This method uses a transformer with multilayer perception for LTR, pre-trained using Lambda loss (Wang et al., 2018) on a different synthetic dataset.
+LR: This method uses a logistic regression-based LTR model, pre-trained on the synthetic dataset.
+NeuralNDCG + TMLP: This method uses a transformer with multilayer perception for LTR, pre-trained using NeuralNDCG loss (Qin et al., 2020) on the synthetic dataset.
+LambdaLoss + TMLP: This method uses a transformer with multilayer perception for LTR, pre-trained using Lambda loss (Wang et al., 2018) on the synthetic dataset.
 5. Experimental Results
@@
 The performance metrics of different models are shown in Table 5.
-The ”Oracle” model has the highest performance across all metrics, which is expected as it serves as the ideal benchmark.
+The ”Oracle” model has the highest performance across all metrics, which is expected as the ideal benchmark.
 The baseline (i.e.
-Google) has the lowest performance, showing that all other models have improved the original ranking.
-The improvements over the baseline are statistically significant for LR, NeuralNDCG + TMLP, and LambdaLoss + TMLP models.
-Despite the simple zero-shot learning nature of the models used in our experiments, the results demonstrate adding verification and reranking into the retrieval process can significantly enhance the performance of a leading baseline search engine.
+Google) has the lowest performance, showing that all other models have improved the original Google ranking when taking execution status into consideration.
+The improvements over the baseline are statistically significant on LR, NeuralNDCG + TMLP, and LambdaLoss + TMLP models.
+Despite their zero-shot learning nature, the results demonstrate adding verification and reranking into the retrieval process can significantly enhance the performance of a leading baseline search engine.
 In comparison, -based rank performs worse than other reranking models.
 This shows reranking using a strong yet single feature is feasible but not good enough.
-Neural ranking models (TMLP) are better than LR probably because they can capture feature interrelationships.
+Two neural ranking models (TMLP) are better than LR probably because they can capture feature interrelationships.
 However, the difference is not very big, possibly due to the low quality of the synthetic data used for pre-training.
@@
 5.2. Further Analysis and Ablation Studies
-Despite the promising performance shown in the zero-shot setting, we want to further explore the correctness based on our data setting and in which cases the system stumbled.
-Therefore, we conducted further analysis in this section to answer several critical questions.
+Despite the promising performance on the zero-shot setting, we want to further explore the correctness based on our data setting and in which cases the system stumbled.
+In this section, we conduct a further analysis to answer several critical questions.
 Q1. When do our proposed methods work and fail? Why?
-To answer this question, an analysis of both success and failure cases are conducted to measure the failure rate at each stage.
+To answer it, success and failure cases have been analyzed to measure the failure rate at each stage.
 After the instruction extraction stage, 398 pages (11.98%) have extracted instructions from the total 3321 pages, with 117 queries (about 70.05%) out of the total 167 queries.
@@
 At this stage, LLMs’ hallucination issue is a major factor: generative content unaligned with original page content induces noises, which decreases the percentage of true positive pages.
-We will further analyze how the extraction performance has been impacted by HTML grounding in question 2.
-At the execution stage, the execution engine filters out 119(29.90%) pages from 68(58.12%) queries and only 279 pages from 91 queries are left.
-For those remaining pages, if the execution engine can finish more than one instruction or the instruction completion degree of the corresponding execution is larger than 0, the extracted instructions are considered as executable and are forwarded to the next stage.
-Among those dropped, 18(15.12%) pages are ones with no action taken by the execution engine and 101 (84.88%)pages have a score of 0 generated by GPT4-V.
-Considering the result of the former is promising, we focus on analysis of : clearly shown in table 6, about 25% of pages with no or partial instructions are set with = 0 and about 70% makes the execution engine struck at the first instruction (the engine can take actions on mobile screen but can’t move forward due to the missing necessary information).
+We will further analyze how HTML grounding has impacted extraction performance in Q2.
+At the execution stage, the execution engine filters 119(29.90%) pages from 68(58.12%) queries, and only 279 pages from 91 queries are left.
+For the remaining pages, if the execution engine can complete more than one instruction or the instruction completion degree of the corresponding execution is greater than 0, the extracted instructions are considered executable and are advanced to the next stage.
+Among those dropped, 18(15.12%) pages have no action taken by the execution proxy, and 101 (84.88%)pages have a score of 0.
+Considering the result of the former is promising, we focus on an analysis based on : clearly shown in table 6, about 25% of pages with no or partial instructions are set with = 0 and about 70% makes the execution engine struck at the first instruction (the engine can’t move forward due to the missing necessary information).
 Another good indicator is that about 75% pages with instructions can support the execution engine to finish more than 75% instructions, which is aligned with human verification.
-However, top 25% pages with no or partial instructions are mistakenly marked as = 1, introducing significant noise in reranking.
-This highlights the necessity of drawing more robust, execution-derived signals from the execution in future work.
+However, top 25% pages with no or partial instructions are mistakenly marked as = 1, introducing significant noise to reranking.
+This highlights the need to incorporate more robust, execution-derived signals in future work.
 | Pages | Mean | Std | Min | 25% | 50% | 75% | Max |
@@
 | No/Partial instructions | 0.3697 | 0.3218 | 0 | 0 | 0.5 | 0.5 | 1 |
-At the reranking stage, we analyze the relationship between features and reranking scores for the best-performed models.
-Most successful cases are because models figure out strong signals among input features, such as and , and learn accurate non-linear weights to reflect on their interrelationship.
-However, current models tend to minimize negative signals even for those with abnormal values: for instance, a feature vector with several very low values, like 2.8e-05, will still be assigned a higher score just because of multiple features that contain higher values.
+In the re-ranking stage, we analyze the relationship between features and reranking scores for the best-performed models.
+Most successful cases occur when models identify strong signals among input features, such as and , and accurately learn non-linear weights to reflect on their interrelationship.
+However, current models tend to minimize negative signals even those with abnormal values: for instance, a feature vector containing several very low values, like 2.8e-05, will still receive a higher score just if it includes multiple features with higher values.
 We attribute this to insufficient negative examples of training and validation sets or limited features incorporated in reranking.
 Q2. What’s the performance of the extraction module? Does grounding matter for extraction? How instruction extract works and how it impacts the final results (failed due to extraction failure):
-We use page URLs collected in ”How-to” WeWeb dataset to fetch the corresponding web pages and apply instruction extraction on individual pages to compare whether the extracted content matches the ground truth steps.
-Accuracy is used to evaluate the instruction extraction performance, which measures whether the model appropriately extracts steps or outputs ”none” based on the given web page and search query.
-GPT4 extraction accuracy is compared with and without step grounding (Table 7).
-The results show IE without grounding (full or partially) is pages out of 169 web pages with relevant instructions.
+We use page URLs collected in the ”How-to” WeWeb dataset to fetch the corresponding web pages and apply instruction extraction on individual pages to compare whether the extracted content matches the ground-truth steps.
+Accuracy is used to evaluate the instruction extraction performance, which measures whether the model appropriately extracts steps or outputs ”none” under different cases.
+Extraction accuracy is listed with and without step grounding (Table 7).
+The results show that instruction extraction without grounding (full or partial) is out of 169 web pages with relevant instructions.
 42 pages with instructions were partially extracted.
 However, it also falsely extracted instructions from 953 pages that don’t contain relevant instructions.
-This happens because the model can hallucinate, extract the wrong set of instructions from a web page with multiple sets of instructions, or significantly paraphrase the instructions.
-With grounding, many partially extracted instructions are corrected by adding missing steps on the same XPaths on the HTML DOM tree, and hallucinated instructions are removed.
-As a result, grounding greatly reduced the number of false positive (i.e.
-extracted instructions from pages without relevant instructions) from 953 to 303, meanwhile reducing the number of true positive (i.e.
-extracted instructions from pages with relevant instructions) from to .
+This happens because the model can hallucinate, extract the wrong instructions from a web page with relevant and irrelevant instructions, or significantly paraphrase the instructions.
+With grounding, many partially extracted instructions are corrected by adding missing steps on the same XPaths on the HTML DOM tree, removing hallucinated instructions.
+As a result, it greatly reduced the number of false positives (i.e., extracted instructions from pages without relevant instructions) from 953 to 303, meanwhile reducing the number of true positives (i.e., extracted instructions from pages with relevant instructions) from to .
 | Extracted Steps | Has relevant instruction | With grounding | Without grounding |
@@
 Q3. Does the zero-shot LTR for reranking help?
-We expect LTR will work if we have training data with the same distribution as the test data.
-However, considering the huge difference between the pre-training and testing data, is it still valid to have a LTR component?
-We tried to rerank by a simple rule: promoting pages with (i.e.
+We expect LTR will work better if we have training data with a similar distribution as the test data.
+However, considering the huge difference between the pre-training and testing data, is it still valid to have an LTR component?
+To clarify this question, we tried to rerank using a simple rule: promoting pages with (i.e.
 GPT4-V predicts the instructions are 100% completed) to the top without changing their related rankings.
-This rule has a , significantly worse than the proposed approach and not significantly better than the baseline.
-This suggests the LTR module is more useful than the simple rule.
-Q4. How sensitive the model performance is for different LTR model settings?
-We varied the TMLP model structure by adding Sigmoid, Tanh activation functions.
-We found the model performance changed a little.
-However, there is nothing statistically significant.
-We expect the small size of our test dataset and the big gap between training and testing data causes the results less sensitive to the perturbations on TMLP.
-The LTR models we trained are far from optimal, and we expect better training data could improve the performance.
+This rule-based reranking reaches , much worse than our proposed LTR approaches while not significantly better than the baseline.
+This suggests LTR models are more useful than a simple rule based on one strong feature.
+Q4. Which features contribute more to the overall performance?
+During the reranking performance analysis, we found that significantly influenced the final ranking: LTR models tend to assign a higher ranking score when values are higher.
+This tendency is more obvious when and are aligned.
+We attribute it to the limited diversity of our synthetic training data, leading to ranking decisions being dominated by strong features.
+Another interesting observation is that the remaining features, which focus on instruction, UI, and action alignments, fail to address some common but tricky circumstances in agent execution, resulting in incorrect ranking decisions.
+For instance, agents may get trapped in ”execution cycles” where robots revert to a previous state due to ambiguous instructions.
+Despite this, the visual features and LTR models still tend to assign higher scores to these instructions.
+This underscores the necessity of incorporating more diverse training data and improving feature representation.
+Q5. How sensitive the model performance is for different LTR model settings?
+We tried Sigmoid, Tanh activation functions in the TMLP models to observe whether a non-linear function will impact the final result.
+Despite a little improvement, there is nothing statistically significant.
+We expect the small size of our test dataset and the big gap between training and testing data to make the results less sensitive to the structural perturbations on TMLP.
+The LTR models we trained are far from optimal, and better training data could improve the performance.
 5.3. Discussion and Future Work
-While our study represents a novel approach for search ranking, we recognize that this is a preliminary study of the problem, and as a result comes with certain limitations.
+Our study represents a novel approach for search reranking.
+However, it is still preliminary and comes with several limitations.
+Therefore, it is necessary to examine those limitations and highlight our future directions.
 Sample size, scope and experimental constraints:
-One of the primary limitations of our study is that it has been done only on the Android platform, and other platforms such as desktop and web browser are not covered and should be studied in the future.
-Although we cover a diverse set of domains, the coverage is not exhaustive.
-Our study also shows the exact performance (P@1 etc.) is domain-specific.
-This may affect the generalization ability of our findings and the reader should take this into account when interpreting the results.
-When applying similar techniques to other platforms or domains, the exact number (P@1 etc.) would be different.
+One limitation is that it has been mainly covered on the Android platform, and other platforms, such as iOS, desktop and web 999We’re now attempting to build a cross-platform execution proxy, following the design mentioned in the section 3.2.3.
+The key idea is to leverage platform-native API to capture accurate control hierarchies and expose UI-level interactions so that the technical concept in this paper can be naturally extended., should be studied in the future, considering their potential difference with Android, especially Accessibility API.
+Although we test diverse mobile application domains, the coverage is not exhaustive.
+Our study also shows that the exact performance (P@1, etc.) is domain-specific.
+This may affect the generalization of our findings, and the reader should consider this when interpreting the results.
+When applying similar techniques to other platforms or domains, the exact number (P@1 etc.) would differ.
 Extension of instruction extraction:
-Although generative LLMs offer a rapid start, the extraction model is significantly hindered by hallucination issues, as noted by Ji et al.
+Although generative LLMs offer a quick start, the hallucination issue has heavily hindered the instruction extraction model, as noted by Ji et al.
 (Ji et al., 2023).
-Our current methodology, which focuses on post-processing with grounding, proves to be a weak regularizer.
-Moving forward, we would like to try generative LMs with multimodal document models such as XDoc (Chen et al., 2022), MarkupLM (Li et al., 2022b), and LayoutLMv3 (Huang et al., 2022b), aiming to enhance in-model grounding efficacy.
+However, during further analysis, we have found that such issue heavily relies on how we represent HTML data to LLMs and how data verification has been organized.
+In our recent experiment, we have improved the algorithm for extracting cleaned HTML and including it in the prompt with a clear XML tag, making LLMs recognize it as a code snippet instead of unstructured text.
+Together with the long context support of gpt4-turbo (GPT-4turbo, 2024), gpt4-o (GPT4o, 2024), a majority of instructions can be extracted successfully in the generative phase.
+In the grounding phase, we have reorganized the context in which the extracted instructions have been verified by the gpt4 models.
+We split the grounding phase into two subphases: the first phase focuses on the relevance between instructions and the given goal, and the next phase reexamizes whether the visual representation of instructions in HTML page is aligned with the goal.
+Through these refinements, we achieved a better result on our current dataset, which can serve as a more promising baseline in our future work.
+Looking forward, we would like to replace those two-phase solutions with a single LLM-extraction model: to be specific, we want to integrate generative LMs with multimodal document models such as XDoc (Chen et al., 2022), MarkupLM (Li et al., 2022b), LayoutLMv3 (Huang et al., 2022b), and LayoutLLM (Luo et al., 2024), aiming to enhance in-model extracting and grounding efficacy.
 Refinement of execution module:
 A more tightly coupled integration of the UI context and the Action Prediction Model can further enhance the execution performance.
-Moreover, our current prompt-based method which used the actionable regions as input does not effectively represent semantic groups on the UI layer.
-For instance, an input box and its label username should be semantically grouped as one region for the execution module.
-Another next step is to enhance the previous work (Hu, 2020), (Liu et al., 2018) to improve the UI contextual information passed to the Action Prediction Model.
+Moreover, our current prompt-based method, which used the actionable regions as input, does not effectively represent semantic groups on the UI layer.
+For example, an input box and its accompanying label, intended for entering a username, should be semantically grouped as a single region.
+Another follow-up is to enhance the previous work (Hu, 2020), (Liu et al., 2018) to improve the UI contextual information passed to the Action Prediction Model.
+In our recent work, by using Dino v1 (Zhang et al., 2022) to understand UI segmentation and injecting into the UI contextual information, the execution module performs slightly better as knowing UI semantic groups.
+However, in the long term, a more promising direction is to follow the line of Ferret v2 (Zhang et al., 2024) by embracing the fine-grained UI representation learnt by Dino v2 (Oquab et al., 2023) into the vision part of the vision language model.
 Reranking model:
-The reranking model we used is very preliminary, using simple feature engineering and no real-world training data.
-With more resources, we should collect more real-world training data with strategic sampling.
-Instead of using manually engineered features, we could represent the verification information, especially the sequence of action screens and instructions, through supervised or unsupervised learning.
-With better reranking models, we would expect the effectiveness of the proposed approach to be more significant.
+Our current reranking model relies on simple feature engineering and does not use real-world training data.
+At the next step, we expect to acquire more real-world training data through strategic sampling.
+Rather than relying on manually engineered features, we could represent verification information, particularly the sequence of action screens and instructions, using more SOTA representation learning techniques.
+Using more realistic data and more advanced representation learning, we anticipate a substantial increase in the effectiveness of our proposed approach.
 Safety issues:
-Although verification agents can save a user time, it might accidentally take harmful actions as some online content could be misleading.
-In practice, we need to implement safeguards to prevent harmful actions from being taken by the execution agent.
+Although verification agents help users avoid tedious manual verification, they might accidentally take harmful actions due to some misleading online content.
+In practice, we need to provide safeguards to prevent these risky actions from being taken by the execution agent.
 Ideally, a client environment that doesn’t impact the user’s real environment should be used.
-If not, we might let each user review and approve extracted instructions or at-risk actions before the automated execution.
-Another approach is to run server-side verification offline and display metadata about verification in the search results (e.g.
-”verified for app1 v3 on android v11”), which allows personalized search results without real time verification on the client side.
+If not, we might allow each user to review and approve extracted instructions or at-risk actions before the automated execution.
+Another approach is to run server-side verification offline and display metadata about verification in the search results (e.g., ”verified for app1 v3 on android v11”), which allows personalized search results without realtime verification on the client side.
 6. Conclusion
-We propose reranking top retrieved results for ”How-to” search queries by promoting candidate web pages with verified executable instructions, which adds an additional layer of usefulness to a typical search engine workflow.
-It ranks search results based on the actual completeness of the candidate instructions, besides traditional metrics such as textual relevance and authority scores.
-The experimental results demonstrate the proposed approach could further improve a very strong baseline search engine(i.e.
+We propose reranking top retrieved results for ”How-to” search queries by promoting candidate web pages with verified executable instructions, which adds an additional layer of usefulness to a traditional search engine workflow.
+It ranks search results based on the actual completeness of candidate instructions, besides traditional metrics such as textual relevance and authority scores.
+The experimental results demonstrate our approach could further improve a very strong baseline search engine(i.e.
 Google).
-This paper is a first step forward in improving the reliability and usefulness of search results of online help resources for ”How-to” queries about software tasks.
-There are several future directions as discussed before, like adapting the system 1 to support verification of other platforms (web, macOS, Windows, iPhone etc.), or collecting user feedback to improve the reranking models and also improving the software agent’s ability to interpret and execute a wider range of instructions.
+This paper is a pioneer work in improving the reliability and usefulness of search results of online help resources for ”How-to” queries about software tasks.
+Along this direction, several promising future research and applications are on the horizon, like adapting the proposed workflow in figure 1 to support verification of other platforms (web, macOS, Windows, iPhone, etc.), or collecting user feedback to improve the reranking models and also improving the software agent’s ability to interpret and execute a wider range of instructions.
 Acknowledgements.
-For ”How-to” WeWeb dataset, we appreciate our research collaborators Sijia Zhong, Valentina Tang, Zoey Zhou, Jimmy Chen, and Chole Wong.
-Without their efforts, it is impossible to finish collecting execution data in a short time period.
-They gave valuable suggestions on MagicWand, making it more mature to support our research work.
-Thank Yaxuan Wang for her efforts to refine figures used in this paper, helping illustrate key ideas concisely and elegantly.
-Annabelle Miin, Krish Pai, Rithvik Chavali, Saket Pathak have involved this research from the beginning and finished work, like deciding on Android Apps, search queries, and guaranteeing data quality via two phases verification.References
+For the ”How-to” WeWeb dataset, we would like to appreciate our research collaborators: Sijia Zhong, Valentina Tang, Zoey Zhou, Jimmy Chen, Hanwen, Yang, and Chole Wong.
+Without their efforts, it would have been impossible to finish data collection on time.
+They also provided valuable suggestions and feedback to MagicWand.
+Additionally, we would like to thank Yaxuan Wang for her effort to refine the figures used to illustrate key ideas concisely and elegantly.
+Finally, we would like to thank our early contributors: Annabelle Miin, Anastasia Miin, Krish Pai, Rithvik Chavali, and Saket Pathak.
+They joined our research at the very beginning and worked tirelessly on data collection and verification.
+Their efforts encompassed sampling search queries, selecting relevant Android apps, and ensuring data accuracy through a rigorous two-stage verification process.References
 - (1)
@@
 - Friedman (2001) Jerome H Friedman. 2001. Greedy function approximation: a gradient boosting machine. Annals of statistics (2001), 1189–1232.
+- GPT-4turbo (2024) GPT-4turbo. 2024. GPT-4turbo. OpenAI. https://platform.openai.com/docs/models/gpt-4-and-gpt-4-turbo
+- GPT4o (2024) GPT4o. 2024. GPT4o. OpenAI. https://platform.openai.com/docs/models/gpt-4o
 - Grover et al. (2019) Aditya Grover, Eric Wang, Aaron Zweig, and Stefano Ermon. 2019. Stochastic optimization of sorting networks via continuous relaxations. arXiv preprint arXiv:1903.08850 (2019).
@@
 569–579.
+- Luo et al.
+(2024) Chuwei Luo, Yufan Shen, Zhaoqing Zhu, Qi Zheng, Zhi Yu, and Cong Yao.
+2024.
+LayoutLLM: Layout Instruction Tuning with Large Language Models for Document Understanding.
+arXiv preprint arXiv:2404.05225 (2024).
 - Mialon et al.
@@
 arXiv:2303.08774 [cs.CL]
+- Oquab et al.
+(2023) Maxime Oquab, Timothée Darcet, Théo Moutakanni, Huy Vo, Marc Szafraniec, Vasil Khalidov, Pierre Fernandez, Daniel Haziza, Francisco Massa, Alaaeldin El-Nouby, et al.
+2023.
+Dinov2: Learning robust visual features without supervision.
+arXiv preprint arXiv:2304.07193 (2023).
 - Pobrotyn et al.
@@
 arXiv preprint arXiv:2205.12386 (2022).
+- Sarkhel et al.
+(2023) Ritesh Sarkhel, Binxuan Huang, Colin Lockard, and Prashant Shiralkar.
+2023.
+Self-training for label-efficient information extraction from semi-structured web-pages.
+Proceedings of the VLDB Endowment 16, 11 (2023), 3098–3110.
+- Sood and Dandapat (2023) Ojasvin Sood and Sandipan Dandapat.
+2023.
+Problematic Webpage Identification: A Trilogy of Hatespeech, Search Engines and GPT.
+In The 7th Workshop on Online Abuse and Harms (WOAH).
+126–137.
 - Sun et al.
@@
 - Zhan and Zhang (2023) Zhuosheng Zhan and Aston Zhang. 2023. You only look at screens: Multimodal chain-of-action agents. arXiv preprint arXiv:2309.11436 (2023).
+- Zhang et al.
+(2022) Hao Zhang, Feng Li, Shilong Liu, Lei Zhang, Hang Su, Jun Zhu, Lionel M Ni, and Heung-Yeung Shum.
+2022.
+Dino: Detr with improved denoising anchor boxes for end-to-end object detection.
+arXiv preprint arXiv:2203.03605 (2022).
+- Zhang et al.
+(2024) Haotian Zhang, Haoxuan You, Philipp Dufter, Bowen Zhang, Chen Chen, Hong-You Chen, Tsu-Jui Fu, William Yang Wang, Shih-Fu Chang, Zhe Gan, et al.
+2024.
+Ferret-v2: An Improved Baseline for Referring and Grounding with Large Language Models.
+arXiv preprint arXiv:2404.07973 (2024).
 - Zhang and Zhang (2023) Zhuosheng Zhang and Aston Zhang. 2023. You Only Look at Screens: Multimodal Chain-of-Action Agents. arXiv e-prints (2023), arXiv–2309.